### 06 Apr 2026 v1

Abstract
Whether polynomial iteration complexity can be established for smoothing Newton methods (SNMs) in symmetric cone programming (SCP) remains a long-standing open problem. A key difficulty lies in the lack of an analogue of the self-concordant convex framework in interior-point methods (IPMs). In this paper, we answer this question affirmatively. We introduce a reduced smoothing barrier augmented Lagrangian (SBAL) function and prove that it is self-concordant convex-concave, which extends the classical self-concordant theory beyond the convex setting. Furthermore, we show that the parameterized smooth equations associated with SNMs are equivalent to the first-order optimality conditions of a minimax problem whose objective is the reduced SBAL function. Motivated by this equivalence, we propose a path-following smoothing Newton method (PFSNM). The reduced SBAL function induces a central path and an associated neighborhood, which provide estimates of the Newton decrement needed for the path-following analysis. As a result, the method is proven to achieve an iteration complexity of $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, matching the best-known short-step bound for IPMs. Numerical results on standard benchmarks show that PFSNM is competitive with several well-known interior-point solvers, providing computational support for the polynomial iteration complexity.
1 Introduction
Symmetric cone programming (SCP) is a fundamental class of convex optimization problems, including linear programming (LP), second-order cone programming (SOCP), semidefinite programming (SDP), and their Cartesian products. Let $\mathbb{E}$ be a Euclidean Jordan algebra equipped with a bilinear operation $\circ$” and an identity element $e$, and let $\mathbb{K}\subseteq\mathbb{E}$ be the associated symmetric cone, i.e., a closed convex cone that is both self-dual and homogeneous. We consider the standard primal-dual form of SCP:
$(\operatorname{P})$ $\min\,\left\{\langle{c},{x}\rangle\,|\,\mathcal{A}{x}={b},\,x\in\mathbb{K}\right\},$ (1)
$(\operatorname{D})$ $\max\left\{\langle{b},\lambda\rangle\,|\,\mathcal{A}^{*}\lambda+{s}={c},\,s\in\mathbb{K}\right\},$
where $c,\,x,\,s\in\mathbb{E}$ and $b,\,\lambda\in\mathbb{R}^{m}$. The linear operator $\mathcal{A}:\mathbb{E}\rightarrow\mathbb{R}^{m}$ is assumed to be surjective, and $\mathcal{A}^{*}$ denotes its adjoint.
Interior-point methods (IPMs) are a standard class of algorithms for symmetric cone programming. They replace the complementarity condition in the Karush–Kuhn–Tucker (KKT) system of (1) by a perturbed relation
$\mathcal{A}x=b,\;\mathcal{A}^{*}\lambda+s=c,\;x,s\in\operatorname{int}\,(\mathbb{K}),\;x\circ s=\mu e,$ (2)
and trace the resulting central path as $\mu\downarrow 0$. Their complexity theory is based on a self-concordant convex framework, which induces a local metric, yields estimates for the Newton decrement, and provides a natural way to define neighborhoods of the central path. In SCP, this framework leads to the classical polynomial iteration bounds: the short-step bound is of order $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$ de Klerk (2002); de Klerk and Vallentin (2016); Schmieta and Alizadeh (2003); Vavasis and Ye (1996); Wright (1997) and the long-step bound is of order $\mathcal{O}(\nu\ln(1/\varepsilon))$ de Klerk (2002); Nesterov (1997); Nocedal and Wright (2006); Schmieta and Alizadeh (2003); Wright (1997), where $\nu$ denotes the rank of $\mathbb{K}$ and $\varepsilon$ is the target accuracy. This worst-case complexity explains why IPMs admit a remarkably robust global theory while retaining strong practical performance.
Alongside IPMs, smoothing Newton methods (SNMs) constitute another important class of algorithms for SCP problems. The basic idea is to reformulate the KKT conditions as a nonsmooth system and then replace the nonsmooth complementarity relation by a parameterized smooth equation
$\mathcal{A}x=b,\;\mathcal{A}^{*}\lambda+s=c,\;\Phi(x,s;\mu)=0,$ (3)
where $\Phi$ is a chosen smoothing function and $\mu$ is driven to zero via a continuation strategy Chen and Tseng (2003); Huang et al. (2004); Kanzow (1996); Peng and Lin (1999). Common choices for $\Phi$ include the smoothing Fischer–Burmeister (FB) function Kanzow (1996); Qi et al. (2000) and the smoothing Chen–Harker–Kanzow–Smale (CHKS) function Chen and Harker (1993); Kanzow (1996); Smale (2000). Unlike IPMs, SNMs do not require the iterates to remain strictly in the interior and are therefore sometimes called non-interior continuation methods. The first non-interior path-following method was proposed by Chen and Harker Chen and Harker (1993). Subsequently, Burke and Xu Burke and Xu (1998) established the first global linear convergence result for a non-interior path-following method for linear complementarity problems, and later proved its local quadratic convergence under suitable assumptions Burke and Xu (2000). Qi, Sun, and Zhou Qi et al. (2000) gave a new formulation of smoothing Newton methods for nonlinear complementarity problems and box-constrained variational inequalities, thereby providing a unified framework that strongly influenced later developments of SNMs. For further advances and related results on SNMs, we refer the reader to Chan and Sun (2008); Huang et al. (2004); Kanzow and Pieper (1999); Kong et al. (2008); Liang et al. (2024); Sun et al. (2004) and the references therein.
Despite these appealing convergence results, SNMs have long lacked the polynomial iteration complexity guarantee, which remains a long-standing open problem Burke and Xu (1998); Kanzow (1996). A key difficulty is that classical SNMs lack an analogue of the self-concordant convex framework that underlies the polynomial complexity theory of IPMs. The framework provides the central path, its neighborhood structure defined by the merit function, and the Newton-decrement estimates required for path-following analysis. These are fundamental, as they give exact estimates for the descent of the merit function when $\mu$ changes, which are typically hard to obtain without the framework.
Several attempts have been made to address this problem. One direction is to integrate parameterized smooth equations into an interior-point framework. A representative example is the interior-point path-following algorithm proposed by Xu and Burke Xu and Burke (1999), which uses the smoothing CHKS function to generate a rescaled Newton direction within the interior-point framework and achieves polynomial bounds. While important, the iterates are still required to stay in the interior of the cone. In contrast, Hotta, Inaba, and Yoshise Hotta et al. (2000) proposed an SNM based on the smoothing CHKS function that eliminates the interior-point requirement, but at the cost of a non-polynomial complexity of $\mathcal{O}\left({\varepsilon^{-6}}\ln{\varepsilon^{-2}}\right)$, which is far from the standard polynomial complexity bounds of IPMs. Hence, the central question is still open:
Can SNMs for SCP attain the polynomial iteration complexity known for IPMs?
This paper answers this question affirmatively. Our starting point is a reduced smoothing barrier augmented Lagrangian (SBAL) function, which reveals the minimax structure hidden in SNMs. We prove that the function is self-concordant convex-concave — a property that extends the classical self-concordant convex framework to saddle-point problems Nemirovski (1999) and admits a global Newton theory for minimax optimization analogous to that for convex minimization. A further key observation is that the parameterized smooth equations associated with SNMs are exactly the first-order optimality conditions of a minimax problem whose objective is the reduced SBAL function. This equivalence induces a local metric for SNMs, enabling the definition of a central path and an associated neighborhood analogous to those in the interior-point framework. More importantly, it provides the Newton decrement estimates required to control the path-following process. Motivated by this equivalence, we propose a path-following smoothing Newton method (PFSNM) for SCP and establish a worst-case iteration complexity of $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, which matches the best-known short-step bound for IPMs on symmetric cones. To the best of our knowledge, this is the first polynomial iteration complexity result for a smoothing Newton method in the general SCP setting.
Although our main focus is theoretical, the resulting method is also computationally attractive. PFSNM admits a Newton system with an explicit Schur complement structure, leading to a more efficient system-formation procedure than in existing SNMs. Furthermore, numerical experiments on standard benchmark problems show that PFSNM is competitive with several well-known interior-point solvers. These results are consistent with the established polynomial-complexity theory.
1.1 Organization
The remainder of the paper is organized as follows. Section 2 reviews preliminaries on Jordan algebras and self-concordant convex-concave functions. Section 3 introduces the reduced SBAL function, establishes its self-concordant convex-concave property, and characterizes the parameterized smooth equations via an equivalent minimax formulation. Section 4 presents the proposed path-following smoothing Newton method and the associated merit functions. Section 5 analyzes the effect of updating the smoothing parameter on the merit functions and finally derives the polynomial iteration complexity of the proposed method. Numerical results are reported in Section 6, validating the effectiveness of PFSNM. The paper concludes in Section 7.
1.2 Notation
[…]
$\langle h_{1},D^{2}f(x)h_{2}\rangle=D^{2}f(x)[h_{1},h_{2}],\quad\forall h_{1},h_{2}\in\mathbb{E}.$
Let $\nabla f(x)$ be the gradient of $f$ at $x$. Then,
$\langle\nabla f(x),h\rangle=Df(x)[h],\quad\forall h\in\mathbb{E}.$
For $g\in C^{k}(\hat{\mathbb{E}}\times\mathbb{E})$, denote by $\nabla_{\hat{x}}g(\hat{x},s)\in\hat{\mathbb{E}}$ the partial gradient of $g$ with respect to $\hat{x}$ at $(\hat{x},s)$, and by $D_{\hat{x}}g(\hat{x},s)$ the corresponding partial derivative of $g$. Let $\mathcal{W},\mathcal{H}:\mathbb{E}\to\mathbb{E}$ be linear operators. Write $\mathcal{H}\succ\mathcal{W}$ if
$\langle h,\mathcal{H}h\rangle>\langle h,\mathcal{W}h\rangle,\quad\forall h\in\mathbb{E}\setminus\{0\}.$
In particular, $\mathcal{H}\succ 0$ means $\langle h,\mathcal{H}h\rangle>0$ for all nonzero $h\in\mathbb{E}$. The boundary and interior of a cone $\mathbb{K}$} are denoted by ${\rm bd}\,(\mathbb{K})$ and ${\rm int}\,({\mathbb{K}})$, respectively. Additional notations and symbols will be introduced as needed.
2 Preliminaries
This section reviews the fundamental concepts that form the basis for our subsequent analysis and algorithmic development. We begin by introducing Euclidean Jordan algebras, which provide the algebraic foundation for symmetric cones and thus play a central role in symmetric cone optimization. We then revisit and generalize the theory of $\alpha$-self-concordant convex-concave functions.
2.1 Euclidean Jordan algebras and symmetric cones
Definition 1
A Euclidean Jordan algebra $(\mathbb{E},\circ,\langle\cdot,\cdot\rangle)$ is a finite-dimensional real inner product space equipped with a bilinear mapping $\circ:\mathbb{E}\times\mathbb{E}\to\mathbb{E}$ such that, for all $x,y\in\mathbb{E}$,
$x\circ y=y\circ x,\ \,x^{2}\circ(x\circ y)=x\circ(x^{2}\circ y),\ \,\langle x\circ y,z\rangle=\langle y,x\circ z\rangle,$ (4)
where $x^{2}:=x\circ x$. The algebra possesses a unique identity element $e$, satisfying $x\circ e=x$ for all $x\in\mathbb{E}$.
A crucial property of Euclidean Jordan algebras is the existence of a spectral decomposition for any element, which generalizes the eigenvalue decomposition of a symmetric matrix.
Theorem 2.1
Let $\mathbb{E}$ be a Euclidean Jordan algebra of rank $\nu$. For any element $z\in\mathbb{E}$, there exist pairwise orthogonal primitive idempotents $\{v_{1},\ldots,v_{\nu}\}$ and unique real eigenvalues $\lambda_{1}(z),\ldots,\lambda_{\nu}(z)$ such that
$z=\sum_{i=1}^{\nu}\lambda_{i}(z)v_{i},$ (5)
where the idempotents satisfy
$\sum_{i=1}^{\nu}v_{i}=e,\ v_{i}\circ v_{j}=0\ \text{for all }i\neq j,\ \text{and}\ v_{i}\circ v_{i}=v_{i}\ \text{for all }i.$
By the spectral decomposition, the determinant of $z$ is defined analogously to that of a real matrix:
$\det(z):=\prod\limits_{i=1}^{\nu}\lambda_{i}(z).$ (6)
An element $z$ lies in the interior of the cone $\mathbb{K}$ if and only if all its eigenvalues are strictly positive; in particular, $\det(z)>0$. The spectral decomposition also enables a functional calculus on $\mathbb{E}$. Given a scalar function $g:\mathbb{R}\to\mathbb{R}$ and an element $z\in\mathbb{E}$ with the spectral decomposition $z=\sum_{i=1}^{\nu}\lambda_{i}(z)v_{i}$, define
$g(z):=\sum_{i=1}^{\nu}g(\lambda_{i}(z))v_{i}.$ (7)
This definition allows for operations such as the square root $z^{1/2}$, the inverse $z^{-1}$, and the logarithm $\ln(z)$, provided that $z\in\operatorname{int}\,(\mathbb{K})$.
In the following, we present three commonly used symmetric cones and their algebraic properties, which are essential to the development of our algorithm.
Table 1: Common types of symmetric cones.
Cone Mathematical representation Jordan product Spectral decomposition
Nonnegative orthant $\mathbb{R}_{+}^{n}=\{x\in\mathbb{R}^{n}\mid x_{i}\geq 0,\,\forall i\}$ $x\circ y:=\operatorname{Diag}(x)y$ $x=\sum_{i=1}^{n}x_{i}e_{i}$
[…]
2.2 Self-concordant convex-concave functions
In this subsection, we introduce the concept and basic properties of $\alpha$-self-concordant convex-concave functions. They provide the theoretical tools for analyzing Newton’s method for finding saddle points, and extend the notion of self-concordant convex functions.
Definition 2((Nemirovski, 1999, Definition 2.1))
Let $\mathbb{E}$ be a Euclidean Jordan algebra, $\mathbb{X}\subset\mathbb{E}$ be an open convex domain, and $\alpha>0$. A convex function $f\in C^{3}(\mathbb{X})$ is called $\alpha$-self-concordant on $\mathbb{X}$ if the following conditions hold:
  1. (i)
    $f$ is a barrier for $\mathbb{X}$, i.e., $f(x^{(k)})\to\infty$ along every sequence of points $x^{(k)}\in\mathbb{X}$ converging to the boundary of $\mathbb{X}$;
  2. (ii)
    For all $x\in\mathbb{X}$ and $h_{x}\in\mathbb{E}$,
    $\left|D^{3}f(x)[h_{x},h_{x},h_{x}]\right|\leq\frac{2}{\alpha^{1/2}}\left(D^{2}f(x)[h_{x},h_{x}]\right)^{3/2}.$ (8)
If $\alpha=1$, $f$ is called standard self-concordant. An $\alpha$-self-concordant convex function $f$ is said to be nondegenerate if the quadratic form $D^{2}f(x)\succ 0$ for all $x\in\mathbb{X}$.
It is well known that an $\alpha$-self-concordant convex function admits a Dikin’s ellipsoid bound, which characterizes the local geometry induced by its second-order derivative.
Theorem 2.2((Nesterov and Nemirovskii, 1994, Theorem 2.1.1))
Let $\mathbb{E}$ be a Euclidean Jordan algebra and $\mathbb{X}\subset\mathbb{E}$ be an open convex domain. Let $f$ be an $\alpha$-self-concordant convex function on $\mathbb{X}$ and $x,\,\Delta x\in\mathbb{E}$. If $r:=\sqrt{\frac{1}{\alpha}D^{2}f(x)[\Delta x,\Delta x]}<1$, then $x+\Delta x\in\mathbb{X}$ and for all $h_{x}\in\mathbb{E}$,
$(1-{r})^{2}D^{2}f(x)[h_{x},h_{x}]\leq D^{2}f(x+\Delta x)[h_{x},h_{x}]\leq\frac{1}{(1-{r})^{2}}D^{2}f(x)[h_{x},h_{x}].$ (9)
Each symmetric cone $\mathbb{K}$ admits a natural barrier function (see (Vieira, 2007, Section 2.6)) $\phi:\operatorname{int}\,(\mathbb{K})\to\mathbb{R}$, defined as
$\phi(x):=-\ln(\det(x)).$ (10)
It follows from Hauser and Güler (2002) that $\phi\in C^{\infty}(\operatorname{int}\,(\mathbb{K}))$ is a nondegenerate 11-self-concordant convex function on $\operatorname{int}\,(\mathbb{K})$. For every $x\in\operatorname{int}\,(\mathbb{K})$, the second-order derivative $D^{2}\phi(x)\succ 0$. Consequently, the inverse operator $(D^{2}\phi(x))^{-1}:\mathbb{E}\to\mathbb{E}$ is well defined and $(D^{2}\phi(x))^{-1}\succ 0$. In particular, the gradient and second-order derivative of $\phi$ satisfy the following properties:
$\nabla\phi(x)=-x^{-1},\quad\langle\nabla\phi(x),(D^{2}\phi(x))^{-1}\nabla\phi(x)\rangle=\nu,\quad\forall x\in\operatorname{int}\,(\mathbb{K}).$ (11)
The subsequent analysis focuses on $\alpha$-self-concordant convex-concave functions. We consider an unconstrained minimax problem whose objective is convex in the minimization variables and concave in the maximization variables. To extend Newton’s method to this setting, it is essential to identify a class of convex-concave functions that exhibit similarly favorable geometric properties of self-concordant convex functions. This motivates the introduction of $\alpha$-self-concordant convex-concave functions, a generalization of the definition proposed in Nemirovski (1999).
Definition 3
Let $\hat{\mathbb{E}}$ and $\mathbb{E}$ be finite-dimensional Euclidean spaces, $f(\hat{x},s)\in C^{3}(\hat{\mathbb{E}}\times\mathbb{E}),$ and $\alpha>0$. The function $f$ is called $\alpha$-self-concordant convex-concave on $\hat{\mathbb{E}}\times\mathbb{E}$ if the following conditions hold:
  1. (i)
    $f$ is convex in $\hat{x}\in\hat{\mathbb{E}}$ for every $s\in\mathbb{E}$, and concave in $s\in\mathbb{E}$ for every $\hat{x}\in\hat{\mathbb{E}}$.
  2. (ii)
    For every $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$,
    $\left|D^{3}f(w)[h,h,h]\right|\leq\dfrac{2}{\alpha^{1/2}}\left(S_{f}(w)[h,h]\right)^{3/2},$ (12)
    where $S_{f}(w)[h,h]:=D^{2}_{\hat{x}\hat{x}}f(w)[h_{\hat{x}},h_{\hat{x}}]-D^{2}_{ss}f(w)[h_{s},h_{s}].$
If $\alpha=1$, the function is called standard self-concordant convex-concave. An $\alpha$-self-concordant convex-concave function $f$ is called nondegenerate if the quadratic form $S_{f}(w)$ is positive definite for all $w\in\hat{\mathbb{E}}\times\mathbb{E}$.
Remark 1
[…]
The following proposition relates nondegenerate $\alpha$-self-concordant convex-concave functions to $\alpha$-self-concordant convex functions.
Proposition 1
Let $\hat{\mathbb{E}}$ and $\mathbb{E}$ be finite-dimensional Euclidean spaces, and let $f(\hat{x},s)$ be a nondegenerate $\alpha$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. Then the following properties hold:
  1. (i)
[…]
  2. (ii)
    For every $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and $h_{1},\,h_{2},\,h_{3}\in\hat{\mathbb{E}}\times\mathbb{E}$, it holds that
    $\left|D^{3}f(w)[h_{1},h_{2},h_{3}]\right|\leq\dfrac{2}{\alpha^{1/2}}\prod_{i=1}^{3}\sqrt{S_{f}(w)[h_{i},h_{i}]}.$
Proof
The proof is provided in Appendix A.1.
Let $f$ be a nondegenerate $\alpha$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. For any vector $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, we define two local norms of $h$ associated with $f$ at $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ by
$\|h\|_{f,w,\alpha}=\sqrt{\frac{1}{\alpha}S_{f}(w)[h,h]},\,\,\|h\|^{*}_{f,w,\alpha}=\sqrt{\frac{1}{\alpha}(S_{f}(w))^{-1}[h,h]}.$ (13)
Since $\alpha$ is intrinsic to $f$, we omit it for brevity and write $\|h\|_{f,w}$ and $\|h\|_{f,w}^{*}$ as shorthand for $\|h\|_{f,w,\alpha}$ and $\|h\|_{f,w,\alpha}^{*}$, respectively.
We introduce three merit functions to measure how far the current iterate is from satisfying the optimality conditions. Specifically, let
$\delta(w):=\|\Delta w\|_{f,w},\,\,\xi(w):=\|\nabla f\|^{*}_{f,w},\,\,\theta(w):=\max\limits_{\tilde{s}}f(\hat{x},\tilde{s})-\min\limits_{\tilde{x}}f(\tilde{x},s),$ (14)
[…]
$\tilde{\delta}_{\hat{x}}(w)=\sqrt{\frac{1}{\alpha}\langle\widetilde{\Delta x},D_{\hat{x}\hat{x}}^{2}f(w)\widetilde{\Delta x}\rangle},\,\,\tilde{\delta}_{s}(w)=\sqrt{-\frac{1}{\alpha}\langle\widetilde{\Delta s},D_{ss}^{2}f(w)\widetilde{\Delta s}\rangle}$, (15)
[…]
Theorem 2.3
Let $\hat{\mathbb{E}}$ and $\mathbb{E}$ be finite-dimensional Euclidean spaces, and let $f:\hat{\mathbb{E}}\times\mathbb{E}\to\mathbb{R}$ be a nondegenerate $\alpha$-self-concordant convex-concave function. For every $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, the following conclusions hold:
  • (i)
    If $r=\|\Delta w\|_{f,w}<1$, then
    $(1-{r})^{2}S_{f}(w)[h,h]\leq S_{f}(w+\Delta w)[h,h]\leq\frac{1}{(1-{r})^{2}}S_{f}(w)[h,h].$ (16)
  • (ii)
    $\delta(w)\leq\xi(w)$.
  • (iii)
    Let $w^{+}:=w+\Delta w\in\hat{\mathbb{E}}\times\mathbb{E}$. If $\delta(w)<1$, then
    $\xi(w^{+})\leq\left(\dfrac{\delta(w)}{1-\delta(w)}\right)^{2}\leq\left(\frac{\xi(w)}{1-\xi(w)}\right)^{2}.$ (17)
    Further, if $\delta(w)\leq\xi(w)\leq 2-\sqrt{3}$, then $\xi(w^{+})\leq\dfrac{\delta(w)}{2}\leq\dfrac{\xi(w)}{2}$.
  • (iv)
    If $w\in K(\theta)$, then
    $\frac{\alpha\xi^{2}(w)}{2(1+\xi(w))}\leq\theta(w).$ (18)
  • (v)
    If $\xi(w)<\frac{1}{3}$, then
    $\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}\leq 1-(1-3\xi(w))^{\frac{1}{3}}.$ (19)
    Further, if $\xi(w)\leq 0.1$, then $\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}<0.2$.
Proof
The proofs of (i)(i)$(iv)$ follow directly from (Nemirovski, 1999, Proposition 2.3 and 5.1). The proof of $(v)$ is provided in Appendix A.2.
3 A minimax reformulation of the smoothing Newton method
In this section, we provide a minimax reformulation of the smoothing Newton method. We first review the classical SNM based on the smoothing CHKS function $\Phi$ for the SCP problem (1).
3.1 From smoothing CHKS function to the reduced SBAL function
Given any point $(x,s)\in\mathbb{E}\times\mathbb{E}$ and $\mu>0$, the classical SNM Engelke and Kanzow (2002); Kanzow and Nagel (2002); Liu et al. (2006) based on $\Phi$ for the SCP (1) (inexactly) solves the parameterized smooth equations
$\mathcal{A}x=b,\;\mathcal{A}^{*}\lambda+s=c,\;\Phi(x,s;\mu)=0.$ (20)
Applying Newton’s method to the parameterized smooth equations (20) yields the linearized system:
$\begin{pmatrix}\mathcal{A}&0&0\\ 0&\mathcal{I}_{\mathbb{E}}&\mathcal{A}^{*}\\ D_{x}\Phi(x,s;\mu)&D_{s}\Phi(x,s;\mu)&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta s\\ \Delta\lambda\end{pmatrix}=-\begin{pmatrix}\mathcal{A}x-b\\ \mathcal{A}^{*}\lambda+s-c\\ \Phi(x,s;\mu)\end{pmatrix}$, (21)
where $\mathcal{I}_{\mathbb{E}}:\mathbb{E}\to\mathbb{E}$ is an identity mapping. The classical SNM proceeds as follows. Starting from $(x^{(0)},s^{(0)},\lambda^{(0)})$, it computes the Newton direction given by (21), performs a line search along this direction at each iteration, and progressively updates the parameter $\mu$ toward zero.
To generalize the parameterized smooth equations, we derive an equivalent characterization of $\Phi$. Let $\phi(x)=-\ln\det(x)$ be the natural barrier of $\mathbb{K}$. Given a fixed $\rho\geq 1$, consider the following optimization problem:
$\min_{z\in\mathrm{int}\,(\mathbb{K})}\Big\{\mu\phi(z)+\langle s,z\rangle+\dfrac{\rho}{2}\|z-x\|^{2}\Big\}.$ (22)
The solution to (22) corresponds to the proximal mapping of the proper closed convex function $\mu\phi(\cdot)+\langle s,\cdot\rangle$ at $x$, which exists uniquely by (Beck, 2017, Theorem 6.3). We denote this solution by $z_{\rho}(x,s;\mu)$. Then
$z_{\rho}(x,s;\mu)=\frac{\rho x-s+\left((\rho x-s)^{2}+4\rho\mu e\right)^{1/2}}{2\rho}\in\operatorname{int}\,(\mathbb{K}),$
and satisfies the optimality condition of (22):
$\mu\nabla\phi(z_{\rho}(x,s;\mu))+s+\rho(z_{\rho}(x,s;\mu)-x)=0.$ (23)
For brevity, we write $z_{\rho}$ instead of $z_{\rho}(x,s;\mu)$ whenever no confusion arises. The natural generalization of $\Phi$ follows as
$\Phi_{\rho}(x,s;\mu):=2(x-z_{\rho}(x,s;\mu)),\quad\forall x,s\in\mathbb{E},\ \mu>0,\ \text{and }\rho\geq 1.$ (24)
In particular, $\Phi(x,s;\mu)=\Phi_{1}(x,s;\mu)$. The parameterized smooth equations (20) generalize to
$\mathcal{A}x=b,\;\mathcal{A}^{*}\lambda+s=c,\;\Phi_{\rho}(x,s;\mu)=0.$ (25)
Applying Newton’s method to (25) yields
$\begin{pmatrix}\mathcal{A}&0&0\\ 0&\mathcal{I}_{\mathbb{E}}&\mathcal{A}^{*}\\ D_{x}\Phi_{\rho}(x,s;\mu)&D_{s}\Phi_{\rho}(x,s;\mu)&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta s\\ \Delta\lambda\end{pmatrix}=-\begin{pmatrix}\mathcal{A}x-b\\ \mathcal{A}^{*}\lambda+s-c\\ \Phi_{\rho}(x,s;\mu)\end{pmatrix}$ (26)
The reformulation (25) extends the standard parameterized smooth equations system (20). Thus, we do not distinguish between the generalized SNM and the classical SNM, and simply refer to both as the SNM. In the following discussion, it suffices to focus on how to reformulate (25).
In line with the idea of building a connection between $\Phi_{\rho}$ and an optimization problem, we next relate (25) to a minimax problem. It is straightforward to verify that (25) is equivalent to
$\mathcal{A}x=b,\;c-s+\frac{\rho}{2}\Phi_{\rho}(x,s;\mu)-\mathcal{A}^{*}\lambda=0,\;\frac{1}{2}\Phi_{\rho}(x,s;\mu)=0.$ (27)
Combining (23), (24), and (27) yields the system
$\mathcal{A}x=b,\;c-s-\rho(z_{\rho}-x)-\mathcal{A}^{*}\lambda=0,\;z_{\rho}-x=0,\;\mu\nabla\phi(z_{\rho})+s+\rho(z_{\rho}-x)=0.$ (28)
This system coincides exactly with the first-order optimality conditions of the following minimax problem, whose objective is the smoothing barrier augmented Lagrangian function $L_{\rho}$:
$\min\limits_{x,z}\max\limits_{s,\lambda}\ \left\{L_{\rho}(x,z,s,\lambda;\mu)\right\},$ (29)
where
$L_{\rho}(x,z,s,\lambda;\mu):=\langle{c},{x}\rangle+\mu\phi(z)-\langle\lambda,\mathcal{A}x-b\rangle+\langle s,z-x\rangle+\dfrac{\rho}{2}\left\|z-x\right\|^{2}.$ (30)
This reformulation reveals that the parameterized smooth equations (25) can be interpreted as the first-order optimality conditions of the minimax problem (29). Consequently, the corresponding theoretical properties of the SNM can be investigated via the SBAL function $L_{\rho}$. However, $L_{\rho}$ is degenerate in the $s$-variable, as $D_{ss}^{2}L_{\rho}(x,z,s,\lambda;\mu)=0$. To overcome this difficulty, we eliminate auxiliary variables and derive a reduced form of $L_{\rho}$.
Since $\mathcal{A}$ is surjective, the linear system $\mathcal{A}x=b$ admits a solution for any $b$. Fix an arbitrary feasible point $\bar{x}$ such that $\mathcal{A}\bar{x}=b$. Let $\hat{\mathbb{E}}$ be a finite-dimensional Euclidean space with ${\rm dim}\,\hat{\mathbb{E}}={\rm dim}\,{\rm ker}\,\mathcal{A}.$ Then, there exists an injective linear operator $\mathcal{B}:\hat{\mathbb{E}}\rightarrow\mathbb{E}$ such that
$\mathcal{A}\mathcal{B}=0,\ \quad\mathcal{B}^{*}\mathcal{B}=\mathcal{I}_{\hat{\mathbb{E}}}$,
Every feasible point $x$ satisfying $\mathcal{A}x=b$ can be written uniquely as
$x=\bar{x}+\mathcal{B}\hat{x},\quad\hat{x}\in\hat{\mathbb{E}}$.
Restricting $L_{\rho}$ to the affine set $\{x\mid Ax=b\}$ eliminates the multiplier $\lambda$. We define $\eta_{\rho}(\cdot,\cdot;\mu):\hat{\mathbb{E}}\times\mathbb{E}\rightarrow\mathbb{R}$ by
$\eta_{\rho}(\hat{x},s;\mu)=\min\limits_{z}\left\{L_{\rho}(\bar{x}+\mathcal{B}\hat{x},z,s,\lambda;\mu)\right\},$ (31)
where the value is independent of $\lambda$ because $A(\bar{x}+\mathcal{B}\hat{x})=b$. Writing $x=\bar{x}+\mathcal{B}\hat{x}$, the unique minimizer of (31) is $z_{\rho}(x,s;\mu)$, and thus
$\eta_{\rho}(\hat{x},s;\mu)=\langle c,x\rangle+\mu\phi(z_{\rho}(x,s;\mu))+\langle s,z_{\rho}(x,s;\mu)-x\rangle+\frac{\rho}{2}\|z_{\rho}(x,s;\mu)-x\|^{2}.$ (32)
We refer to $\eta_{\rho}$ as the reduced SBAL function. The corresponding minimax problem is
$\min\limits_{\hat{x}\in\hat{\mathbb{E}}}\max\limits_{s\in\mathbb{E}}\left\{\eta_{\rho}(\hat{x},s;\mu)\right\}.$ (33)
Compared with $L_{\rho}$, the reduced SBAL function $\eta_{\rho}$ admits more favorable structural properties, as established in the following subsection. In the remainder of the paper, we focus exclusively on the minimax problem (33) and $\eta_{\rho}$.
3.2 Properties of the reduced SBAL function
In this subsection, we discuss the properties of the reduced SBAL function $\eta_{\rho}$. Recall that $z_{\rho}(x,s;\mu)$ satisfies the nonlinear equation:
$\mu\nabla\phi(z_{\rho}(x,s;\mu))+s+\rho(z_{\rho}(x,s;\mu)-x)=0.$ (34)
Define the adjoint variable associated with $z_{\rho}(x,s;\mu)$ by
$y_{\rho}(x,s;\mu)=s+\rho(z_{\rho}(x,s;\mu)-x).$
The variables $z_{\rho}(x,s;\mu)$ and $y_{\rho}(x,s;\mu)$ satisfy the following properties.
Lemma 1
For any scalars $\mu>0$ and $\rho\geq 1$, the following statements are equivalent:
$\text{(i)}\;z_{\rho}(x,s;\mu)=x,\,\text{(ii)}\;y_{\rho}(x,s;\mu)=s,\,\text{(iii)}\;x,\,s\in\operatorname{int}\,(\mathbb{K}),\,x\circ s=\mu e.$ (35)
Proof
The equivalence $(i)\Longleftrightarrow(ii)$ follows directly from the definition of $y_{\rho}(x,s;\mu)$. It suffices to prove that $(i)\Longleftrightarrow(iii)$.
Suppose that condition $(iii)$ holds. It follows from (34) that
$z_{\rho}(x,s;\mu)=\frac{\rho x-s+\left((s-\rho x)^{2}+4\rho\mu e\right)^{1/2}}{2\rho}.\\$ (36)
Since $x\circ s=\mu e$,
$z_{\rho}(x,s;\mu)$ $=\frac{\rho x-s+\left(s^{2}-2\rho s\circ x+\rho^{2}x^{2}+4\rho\mu e\right)^{1/2}}{2\rho}=x,$ (37)
which establishes condition $(i)$.
Conversely, assume that condition $(i)$ holds. By (34) and the inclusions $z_{\rho},x\in\operatorname{int}\,(\mathbb{K})$,
$\mu\nabla\phi(x)+s=0.$ (38)
Combining (11) and (38) yields $s\in\operatorname{int}\,(\mathbb{K})$ and $x\circ s=\mu e$. This establishes condition $(iii)$ and completes the proof.
Define the linear operators
${\mathcal{W}}=\dfrac{\mu}{\rho}D^{2}\phi(z_{\rho}(x,s;\mu)),\quad{\mathcal{H}}=\mathcal{I}_{\mathbb{E}}+\mathcal{W}.$ (39)
Since the natural barrier $\phi$ is strictly convex on $\operatorname{int}\,(\mathbb{K})$, we have
$\mathcal{H}\succ\mathcal{W}\succ 0,\quad\mathcal{H}\succ\mathcal{I}_{\mathbb{E}}.$
For brevity, all derivatives with respect to $\mu$ are denoted by a prime. For example, $z_{\rho}^{\prime}(x,s;\mu)=D_{\mu}z_{\rho}(x,s;\mu)$. When no ambiguity arises, we also abbreviate $y_{\rho}(x,s;\mu)$ as $y_{\rho}$. The following theorem characterizes the derivatives of $z_{\rho}(x,s;\mu)$ and $y_{\rho}(x,s;\mu)$.
Theorem 3.1
For any $\rho\geq 1$, the mappings $z_{\rho}(x,s;\mu)$ and $y_{\rho}(x,s;\mu)$ are smooth with respect to $(x,s,\mu)$ on $\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++}$. Moreover, their partial derivatives with respect to $(x,s)$ are given by
[…]
and the derivatives with respect to $\mu$ are given by
$z^{\prime}_{\rho}(x,s;\mu)=-\rho^{-1}\mathcal{H}^{-1}\nabla\phi(z_{\rho}(x,s;\mu)),$ (41)
$y^{\prime}_{\rho}(x,s;\mu)=-\mathcal{H}^{-1}\nabla\phi(z_{\rho}(x,s;\mu)).$
Proof
Recall that $\phi\in C^{\infty}(\operatorname{int}\,(\mathbb{K}))$ and that $z_{\rho}(x,s;\mu)$ is defined as the unique solution to (34). Since
$\mu D^{2}\phi(z_{\rho}(x,s;\mu))+\rho\mathcal{I}_{\mathbb{E}}=\rho\mathcal{H}\succ 0,\quad\forall\,(x,s,\mu)\in\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++},$ (42)
the Jacobian of (34) with respect to $z_{\rho}$ is nonsingular. Hence, the implicit function theorem implies $z\in C^{\infty}(\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++},\mathbb{E})$. By definition,
$y_{\rho}(x,s;\mu)=s+\rho(z_{\rho}(x,s;\mu)-x),$ (43)
which implies that $y_{\rho}\in C^{\infty}(\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++},\mathbb{E})$. Moreover, $y_{\rho}$ satisfies
$\mu\nabla\phi(z_{\rho}(x,s;\mu))+y_{\rho}(x,s;\mu)=0.$ (44)
Differentiating both sides of (43) and (44) with respect to $x$ yields
$D_{x}y_{\rho}(x,s;\mu)=\rho(D_{x}z_{\rho}(x,s;\mu)-\mathcal{I}_{\mathbb{E}})$
and
$\mu D^{2}\phi(z_{\rho}(x,s;\mu))D_{x}z_{\rho}(x,s;\mu)+D_{x}y_{\rho}(x,s;\mu)=0.$
Consequently, we have
$D_{x}z_{\rho}(x,s;\mu)=\rho(\mu D^{2}\phi(z_{\rho}(x,s;\mu))+\rho\mathcal{I}_{\mathbb{E}})^{-1}=\mathcal{H}^{-1}$
and
$D_{x}y_{\rho}(x,s;\mu)$ $=-\rho\left(\dfrac{\mu}{\rho}D^{2}\phi(z_{\rho}(x,s;\mu))+\mathcal{I}_{\mathbb{E}}\right)^{-1}\left(\dfrac{\mu}{\rho}D^{2}\phi(z_{\rho}(x,s;\mu))\right)$
$=-\rho\mathcal{H}^{-1}\mathcal{W}.$
The remaining identities can be proved in the same way.
By Theorem 3.1 and (26), the search direction generated by the SNM is equivalently written as the solution to the linear system
$\begin{pmatrix}\mathcal{A}&0&0\\ 0&\mathcal{I}_{\mathbb{E}}&\mathcal{A}^{*}\\ \mathcal{H}^{-1}\mathcal{W}&\rho^{-1}\mathcal{H}^{-1}&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta s\\ \Delta\lambda\end{pmatrix}=-\begin{pmatrix}\mathcal{A}x-b\\ \mathcal{A}^{*}\lambda+s-c\\ x-z_{\rho}\end{pmatrix}$. (45)
The following proposition guarantees the uniqueness of the above search direction.
Proposition 2
For any point $(x,s,\lambda)$ and scalars $\mu>0$, $\rho\geq 1$, the Newton system (45) generated by the SNM admits a unique solution.
Proof
The third equation in (45) gives
$\Delta s=-\rho\mathcal{W}\Delta x-\rho\mathcal{H}(x-z_{\rho}).$ (46)
Substitute this expression into the remaining equations. It suffices to verify the uniqueness of the solution to
$\begin{pmatrix}-\rho\mathcal{W}&\mathcal{A}^{*}\\ \mathcal{A}&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta\lambda\end{pmatrix}=-\begin{pmatrix}\rho\mathcal{H}(z_{\rho}-x)+\mathcal{A}^{*}\lambda+s-c\\ \mathcal{A}x-b\end{pmatrix}$. (47)
The Schur complement of $\begin{pmatrix}-\rho\mathcal{W}&\mathcal{A}^{*}\\ \mathcal{A}&0\end{pmatrix}$ relative to $\rho\mathcal{W}$ is $\rho^{-1}\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}$ since $\rho\mathcal{W}$ is invertible. Therefore, the system (47) admits a unique solution. This completes the proof.
The next corollary provides explicit expressions for the first- and second-order derivatives of $\eta_{\rho}(\hat{x},s;\mu)$ with respect to $\hat{x}$ and $s$.
Corollary 1
For any $\rho\geq 1$, the reduced SBAL function $\eta_{\rho}(\hat{x},s;\mu)$ is smooth on $\hat{\mathbb{E}}\times\mathbb{E}\times\mathbb{R}_{++}$. Furthermore, for $x=\bar{x}+\mathcal{B}\hat{x}$,
$\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)=\mathcal{B}^{*}(c-y_{\rho}(x,s;\mu)),\,\nabla_{s}\eta_{\rho}(\hat{x},s;\mu)=z_{\rho}(x,s;\mu)-x,$ (48)
and
$D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)=\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B},\quad D^{2}_{ss}\eta_{\rho}(\hat{x},s;\mu)=-\rho^{-1}\mathcal{H}^{-1}.$ (49)
Proof
The smoothness of $\eta_{\rho}$ follows immediately from Theorem 3.1. By differentiating (31) with respect to $x$, we obtain from the equality $\hat{x}=\mathcal{B}^{*}(x-\bar{x})$ and the chain rule that
$\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)$ $=\mathcal{B}^{*}(\nabla_{x}\eta_{\rho}(\hat{x},s;\mu))$ (50)
$=\mathcal{B}^{*}\big(c+\mu D_{x}z_{\rho}(x,s;\mu)\nabla\phi(z_{\rho}(x,s;\mu))$
$\qquad+((D_{x}z_{\rho}(x,s;\mu)-\mathcal{I}_{\mathbb{E}})(s+\rho(z_{\rho}(x,s;\mu)-x))\big)$
$\overset{(i)}{=}\mathcal{B}^{*}\big(D_{x}z_{\rho}(x,s;\mu)(\mu\nabla\phi(z_{\rho}(x,s;\mu))+y_{\rho}(x,s;\mu))$
$\qquad+c-y_{\rho}(x,s;\mu)\big)$
$=\mathcal{B}^{*}(c-y_{\rho}(x,s;\mu)),$
where $(i)$ follows from the definition of $y_{\rho}(x,s;\mu)$. Further differentiating (50) with respect to $\hat{x}$ yields
$D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)=\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}.$ (51)
The remaining identities follow by analogous arguments.
Theorem 3.2
For any $\mu>0$ and $\rho\geq 1$, the reduced SBAL function $\eta_{\rho}(\cdot,\cdot;\mu)$ is a nondegenerate $\mu$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. Furthermore, $\eta_{\rho}(\cdot,s;\mu)$ is nondegenerate $\mu$-self-concordant on $\hat{\mathbb{E}}$ for every $s\in\mathbb{E}$, and $\eta_{\rho}(\hat{x},\cdot;\mu)$ is nondegenerate $\mu$-self-concordant on $\mathbb{E}$ for every $\hat{x}\in\hat{\mathbb{E}}$.
Proof
The smoothness of $\eta_{\rho}$ follows from Corollary 1. For any $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$,
$S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]=\rho\langle h_{\hat{x}},\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}h_{\hat{x}}\rangle+\rho^{-1}\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle.$ (52)
Since $\mathcal{B}^{*}$ is surjective and $\mathcal{W}\succ 0$, $S_{\eta_{\rho}}(\hat{x},s;\mu)$ is positive definite. It suffices to prove that
$|D^{3}\eta_{\rho}(\hat{x},s;\mu)[h,h,h]|\leq\dfrac{2}{\sqrt{\mu}}(S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h])^{3/2}.$ (53)
Let $\hat{h}=(\hat{h}_{\hat{x}},\hat{h}_{s})=(\mathcal{H}^{-1}\mathcal{B}h_{\hat{x}},\mathcal{H}^{-1}h_{s})\in\mathbb{E}\times\mathbb{E}$. By Corollary 1 and the formula $\mathcal{H}=\mathcal{I}_{\mathbb{E}}+\mathcal{W}=\mathcal{I}_{\mathbb{E}}+\frac{\mu}{\rho}D^{2}\phi(z_{\rho})$, we have
$S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]$ (54)
$=$ $\ \rho\langle\hat{h}_{\hat{x}},\mathcal{H}\mathcal{W}\hat{h}_{\hat{x}}\rangle+\rho^{-1}\langle\hat{h}_{s},\mathcal{H}\hat{h}_{s}\rangle$
$=$ $\ \mu D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}},\hat{h}_{\hat{x}}]+\frac{\mu^{2}}{\rho}\|D^{2}\phi(z_{\rho})\hat{h}_{\hat{x}}\|^{2}+\rho^{-1}\|\hat{h}_{s}\|^{2}+\frac{\mu}{\rho^{2}}D^{2}\phi(z_{\rho})[\hat{h}_{s},\hat{h}_{s}]$
$=$ $\ \mu D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s},\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s}]+\dfrac{1}{\rho}\|\mu D^{2}\phi(z_{\rho})\hat{h}_{\hat{x}}+\hat{h}_{s}\|^{2}$
$\geq$ $\ \mu D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s},\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s}],$
and
$D^{2}\eta_{\rho}(\hat{x},s;\mu)[h,h]$ $=\rho\langle\mathcal{B}h_{\hat{x}},(\mathcal{I}_{\mathbb{E}}-\mathcal{H}^{-1})\mathcal{B}h_{\hat{x}}\rangle-\rho^{-1}\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle$ (55)
$\qquad\qquad+2\langle\mathcal{B}h_{\hat{x}},(\mathcal{H}^{-1}-\mathcal{I}_{\mathbb{E}})h_{s}\rangle.$
Direct computation gives
$D_{\hat{x}}\mathcal{H}[h_{\hat{x}},\cdot,\cdot]=\frac{\mu}{\rho}D^{3}\phi(z_{\rho})[\hat{h}_{\hat{x}}$$\cdot,\cdot$], (56a)
$D_{s}\mathcal{H}[h_{s},\cdot,\cdot]=-\frac{\mu}{\rho^{2}}D^{3}\phi(z_{\rho})[\hat{h}_{s},\cdot,\cdot],$ (56b)
$D_{\hat{x}}\mathcal{H}^{-1}[h_{\hat{x}},\cdot,\cdot]=-\frac{\mu}{\rho}\mathcal{H}^{-1}D^{3}\phi(z_{\rho})[\hat{h}_{\hat{x}},\cdot,\cdot]\mathcal{H}^{-1},$ (56c)
$D_{s}\mathcal{H}^{-1}[h_{s},\cdot,\cdot]=\frac{\mu}{\rho^{2}}\mathcal{H}^{-1}D^{3}\phi(z_{\rho})[\hat{h}_{s},\cdot,\cdot]\mathcal{H}^{-1}.$ (56d)
Thus, we have
$D^{3}\eta_{\rho}(\hat{x},s;\mu)[h,h,h]$ (57)
$=$ $\ -\rho D_{\hat{x}}\mathcal{H}^{-1}[h_{\hat{x}},\mathcal{B}h_{\hat{x}},\mathcal{B}h_{\hat{x}}]-\rho D_{s}\mathcal{H}^{-1}[h_{s},h_{\hat{x}},h_{\hat{x}}$]
$\ \qquad-\rho^{-1}D_{\hat{x}}\mathcal{H}^{-1}[h_{\hat{x}},h_{s},h_{s}]-\rho^{-1}D_{s}\mathcal{H}^{-1}[h_{s},h_{s},h_{s}]$
$\ \qquad+2D_{\hat{x}}\mathcal{H}^{-1}[h_{\hat{x}},\mathcal{B}h_{\hat{x}},h_{s}]+2D_{s}\mathcal{H}^{-1}[h_{s},\mathcal{B}h_{\hat{x}},h_{s}]$
$=$ $\ \mu D^{3}\phi(z_{\rho})[\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s},{\hat{h}_{\hat{x}}}-\rho^{-1}\hat{h}_{s},\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s}]$
$\overset{(i)}{\leq}$ ${2\mu}\left\{D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s},\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s}]\right\}^{3/2}$
$\overset{(ii)}{\leq}$ $\ \dfrac{2}{\sqrt{\mu}}\left(S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]\right)^{3/2},$
where $(i)$ follows from the self-concordant property of $\phi$, and $(ii)$ follows from (54). Hence $\eta_{\rho}(\cdot,\cdot;\mu)$ is a nondegenerate $\mu$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. The remaining statements follow immediately from Proposition 1.
3.3 Equivalent characterization of the parameterized smooth equations
This subsection is central to our analysis. We establish the equivalence between the parameterized smooth equations (25) and the first-order optimality conditions of the minimax problem (33).
We begin by recalling the barrier subproblem arising in IPMs:
$\min\left\{\langle{c},{x}\rangle+\mu\phi(x)\,|\,\mathcal{A}x=b,\,x\in\operatorname{int}\,(\mathbb{K})\right\}.$ (58)
The following lemma establishes the connection between (25) and (58).
Lemma 2
For any $\mu>0$ and $\rho\geq 1$, the triple $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth equations (25) if and only if it is a primal-dual optimal solution of the barrier subproblem (58).
Proof
By definition, the triple $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth equations (25) if and only if it satisfies
$\mathcal{A}x(\mu)-b=0,\;\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=0,\;\Phi_{\rho}(x(\mu),s(\mu);\mu)=0.$ (59)
Since $\Phi_{\rho}(x(\mu),s(\mu);\mu)=2(x(\mu)-z_{\rho}(x(\mu),s(\mu);\mu))$, it follows from Lemma 1 that
$\mathcal{A}x(\mu)=b,\,\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=0,\,x(\mu),\,s(\mu)\in\operatorname{int}\,(\mathbb{K}),\,x(\mu)\circ s(\mu)$ $=\mu e.$ (60)
Consequently, $(x(\mu),s(\mu),\lambda(\mu))$ is a primal-dual optimal solution of the barrier subproblem (58). The converse implication follows by reversing the above arguments, and the proof is complete.
Based on Lemma 2, we now establish the equivalence between the parameterized smooth equations (25) and the first-order optimality conditions of the minimax problem (33).
Theorem 3.3
For any $\mu>0$ and $\rho\geq 1$, suppose that $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth equations (25). Then the pair $(\hat{x}(\mu),s(\mu)):=(\mathcal{B}^{*}(x(\mu)-\bar{x}),s(\mu))$ is a saddle point of the minimax problem (33). Conversely, if $(\hat{x}(\mu),s(\mu))$ is a saddle point of (33), then
$(x(\mu),s(\mu),\lambda(\mu)):=(\bar{x}+\mathcal{B}\hat{x}(\mu),s(\mu),(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A}(c-s(\mu)))$
solves the parameterized smooth equations (25).
Proof
Suppose that the triple $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth equations (25). Then by Lemma 2, we have
$\mathcal{A}x(\mu)=b,\,\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=0,\,x(\mu),\,s(\mu)\in\operatorname{int}\,(\mathbb{K}),\,x(\mu)\circ s(\mu)$ $=\mu e.$ (61)
This combined with Lemma 1 yields
$z_{\rho}(x(\mu),s(\mu);\mu)=x(\mu).$ (62)
Thus, it follows from Corollary 1 that
$\nabla_{s}\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu)=z_{\rho}(x(\mu),s(\mu);\mu)-x(\mu)=0.$ (63)
Moreover,
$\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)$ $=\mathcal{B}^{*}(c-y_{\rho}(x(\mu),s(\mu);\mu))$ (64)
$=\mathcal{B}^{*}(c-s(\mu)-\rho(z_{\rho}(x(\mu),s(\mu);\mu)-x(\mu)))$
$\overset{(i)}{=}\mathcal{B}^{*}(c-s(\mu))$
$\overset{(ii)}{=}\mathcal{B}^{*}\mathcal{A}^{*}\lambda(\mu)$
$=0,$
where $(i)$ follows from $z_{\rho}(x(\mu),s(\mu);\mu)=x(\mu)$, and $(ii)$ follows from the second equation in (61). Therefore, $(\hat{x}(\mu),s(\mu)):=(\mathcal{B}^{*}(x(\mu)-\bar{x}),s(\mu))$ satisfies the first-order optimality conditions of the minimax problem (33). Since $\eta_{\rho}$ is convex-concave, this pair is a saddle point of (33).
Conversely, suppose that $(\hat{x}(\mu),s(\mu))$ is a saddle point of (33). Then it satisfies
$z_{\rho}(x(\mu),s(\mu);\mu)=x(\mu),\,\mathcal{B}^{*}(c-y_{\rho}(x(\mu),s(\mu);\mu))=0,$ (65)
where $x(\mu)=\bar{x}+\mathcal{B}\hat{x}(\mu)$. By Lemma 1, we have
$x(\mu),\,s(\mu)\in\operatorname{int}\,(\mathbb{K}),\;x(\mu)\circ s(\mu)=\mu e,$
and $\mathcal{A}x(\mu)-b=\mathcal{A}\bar{x}-b+\mathcal{A}\mathcal{B}\hat{x}(\mu)=0$. It remains to show that
$\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=0.$ (66)
Note that the second equation of (65) implies
$c-y_{\rho}(x(\mu),s(\mu);\mu)=c-s(\mu)\in({\rm ker}\,\mathcal{A})^{\perp}.$
Since $\mathcal{A}$ is surjective, and $\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A}$ is the orthogonal projector onto $({\rm ker}\,\mathcal{A})^{\perp}$, we obtain
$\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=(\mathcal{I}_{\mathbb{E}}-\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A})(s(\mu)-c)=0,$ (67)
which concludes the proof.
Remark 2
According to (Nesterov and Todd, 1998, Theorem 4.1), the barrier subproblem (58) admits a unique primal-dual optimal solution. This combined with Theorem 3.3 implies that the minimax problem (33) admits a unique saddle point for any $\mu>0$ and $\rho\geq 1$.
The following theorem characterizes the search direction of the SNM.
Theorem 3.4
Let $(x,s,\lambda)$ satisfy the linear constraint $\mathcal{A}x=b$. For any scalars $\mu>0$ and $\rho\geq 1$, suppose that $(\Delta x,\Delta s,\Delta\lambda)$ is the search direction generated by the SNM, satisfying (45). Then the pair $(\Delta\hat{x},\Delta s):=(\mathcal{B}^{*}\Delta x,\Delta s)$ is the unique Newton direction for the minimax problem (33), given by
$\begin{pmatrix}\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}&-\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\\ -\mathcal{H}^{-1}\mathcal{W}\mathcal{B}&-\rho^{-1}\mathcal{H}^{-1}\end{pmatrix}\begin{pmatrix}\Delta\hat{x}\\ \Delta s\end{pmatrix}=-\begin{pmatrix}\mathcal{B}^{*}(c-y_{\rho})\\ z_{\rho}-x\end{pmatrix}$. (68)
Conversely, suppose that $(\Delta\hat{x},\Delta s)$ is the Newton direction for the minimax problem (33). Define
$\Delta\lambda:=\left({\mathcal{A}\mathcal{A}^{*}}\right)^{-1}\mathcal{A}(-s-\Delta s+c-\mathcal{A}^{*}\lambda).$ (69)
Then $(\Delta x,\Delta s,\Delta\lambda)=(\mathcal{B}\Delta\hat{x},\Delta s,\Delta\lambda)$ is the unique search direction given by the SNM.
Proof
Suppose that $(\Delta x,\Delta s,\Delta\lambda)$ is the search direction generated by the SNM. Let $(\Delta\hat{x},\Delta s)=(\mathcal{B}^{*}\Delta x,\Delta s)$. Then,
$\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}\Delta\hat{x}-\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\Delta s+\mathcal{B}^{*}(c-y_{\rho})$ (70)
$\overset{(i)}{=}$ $\ \mathcal{B}^{*}(\rho\mathcal{H}^{-1}\mathcal{W}\Delta x-\mathcal{H}^{-1}\mathcal{W}\Delta s+c-s-\rho(z_{\rho}-x))$
$\overset{(ii)}{=}$ $\ \mathcal{B}^{*}(c-s-\Delta s)$
$\overset{(iii)}{=}$ $\ \mathcal{B}^{*}\mathcal{A}^{*}(\lambda+\Delta\lambda)$
$=$ $0,$
where $(i)$ follows from the definition of $y_{\rho}(x,s;\mu)$, $(ii)$ and $(iii)$ are the third and second equations of (45), respectively. The second equation of (68) follows directly from the third equation of (45). Consequently, the equations (68) admit at least one solution. Since the Schur complement of the operator in (68) relative to $\rho^{-1}\mathcal{H}^{-1}$ is $\rho\mathcal{B}^{*}\mathcal{W}\mathcal{B}\succ 0$, the uniqueness follows directly.
Conversely, suppose that $(\Delta\hat{x},\Delta s)$ is the Newton direction for the minimax problem (33). Then,
$\mathcal{A}\Delta x=\mathcal{A}\mathcal{B}\Delta\hat{x}=0=-(\mathcal{A}x-b),$ (71)
which satisfies the first equation of (45). It suffices to verify the second equation. Left-multiplying the second equation in (68) by $\rho\mathcal{B}^{*}$ and adding it to the first equation, we obtain
$\mathcal{B}^{*}(c-s-\Delta s)=0.$ (72)
This implies $s+\Delta s-c\in({\rm ker}\,\mathcal{A})^{\perp}$. Combined with (69),
$\mathcal{A}^{*}(\lambda+\Delta\lambda)+s+\Delta s-c$ $=(\mathcal{I}_{\mathbb{E}}-\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A})(s+\Delta s-c).$ (73)
Since $\mathcal{I}_{\mathbb{E}}-\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A}$ is the orthogonal projection onto ${\rm ker}\,\mathcal{A}$, we have
$\mathcal{A}^{*}(\lambda+\Delta\lambda)+s+\Delta s-c=0,$
which is the second equation of (45). The uniqueness of $(\Delta x,\Delta s,\Delta\lambda)$ follows from Proposition 2, and the proof is complete.
Theorems 3.3 and 3.4 provide an equivalent characterization of both the parameterized smooth equations and the search direction generated by the SNM under consideration via the reduced SBAL function $\eta_{\rho}$. This equivalence implies that, in the subsequent algorithmic analysis, it suffices to study the Newton iterations applied to the minimax problem (33). This offers a convenient and powerful tool for analyzing the behavior of the SNM. We conclude this section with a summary of the main characterizations obtained.
[…]
Figure 1: Summary of equivalence relationships
4 A path-following smoothing Newton method
This section proposes a path-following smoothing Newton method for symmetric cone programming. The method adopts a two-phase structure. In the first phase, an initial point within a well-defined neighborhood of the central path is efficiently constructed. Starting from this point, the second phase iteratively refines a solution within the neighborhood.
4.1 Neighborhood of the central path
For practical implementation and theoretical analysis, maintaining iterates within a well-defined neighborhood of the central path is critical. To measure the proximity of a point to the central path and drive the iteration process, we introduce some important functions.
By Theorem 3.2, the function $\eta_{\rho}(\hat{x},s;\mu)$ is strictly convex in $\hat{x}\in\hat{\mathbb{E}}$ and strictly concave in $s\in\mathbb{E}$. Accordingly, we measure its sub-optimality by the primal-dual gap function following (14):
$\theta_{\rho}(\hat{x},s;\mu)$ $=\max_{\tilde{s}}\eta_{\rho}(\hat{x},\tilde{s};\mu)-\min_{\tilde{x}}\eta_{\rho}(\tilde{x},s;\mu)$ (74)
$=\eta_{\rho}(\hat{x},{s_{\rho}(\hat{x},\mu)};\mu)-\eta_{\rho}(\hat{x}_{\rho}(s,\mu),s;\mu),$
where
$s_{\rho}(\hat{x},\mu)={{\operatorname*{arg\,max}_{\tilde{s}}}\,\eta_{\rho}(\hat{x},\tilde{s};\mu)},\ \hat{x}_{\rho}(s,\mu)={{\operatorname*{arg\,min}_{\tilde{x}}}\,\eta_{\rho}(\tilde{x},s;\mu)}.$
Let $(\hat{x}(\mu),s(\mu))$ be the saddle point of the minimax problem (33), which always exists for any $\mu>0$ and $\rho\geq 1$ by Remark 2. Then for any $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$, the following inequalities hold:
$\max_{\tilde{s}}\eta_{\rho}(\hat{x},\tilde{s};\mu)\geq\min\limits_{\tilde{x}}\max_{\tilde{s}}\eta_{\rho}(\tilde{x},\tilde{s};\mu)=\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu),$ (75)
$\min_{\tilde{x}}\eta_{\rho}(\tilde{x},s;\mu)\leq\max\limits_{\tilde{s}}\min\limits_{\tilde{x}}\eta_{\rho}(\tilde{x},\tilde{s};\mu)=\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu).$
Let ${\rm val}\,(\mathrm{P}_{\mu})$ denote the optimal value of the primal barrier problem (58). By Theorem 3.3,
$\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu)={\rm val}\,(\mathrm{P}_{\mu}).$
This combined with (75) implies that, for any $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$|\eta_{\rho}(\hat{x},s;\mu)-{\rm val}\,(\mathrm{P}_{\mu})|\leq\theta_{\rho}(\hat{x},s;\mu).$ (76)
Moreover, (75) implies that $\theta_{\rho}(\hat{x},s;\mu)\geq 0$. The equality holds if and only if $(\hat{x},s)$ is the saddle point of problem (33). This gap therefore provides a rigorous certificate of optimality and will be used to measure the proximity to the central path.
In practice, evaluating the exact primal-dual gap $\theta_{\rho}(\hat{x},s;\mu)$ is computationally prohibitive, as it requires solving optimization problems. Recall that $x=\bar{x}+\mathcal{B}\hat{x}$. Let $(\Delta x,\Delta s,\Delta\lambda)$ be the search direction given by (45), and let $\Delta w=(\Delta\hat{x},\Delta s)$ be the Newton direction for the minimax problem (33). Define $w:=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$. For algorithmic purposes, we follow (14) and introduce easily computable merit functions:
$\delta_{\hat{x},\rho}(\hat{x},s;\mu)$ $=\sqrt{\dfrac{1}{\mu}\langle\Delta\hat{x},{D_{\hat{x}\hat{x}}^{2}{\eta_{\rho}(\hat{x},s;\mu)}}\Delta\hat{x}\rangle}$, (77a)
$\delta_{s,\rho}(\hat{x},s;\mu)$ $=\sqrt{-\dfrac{1}{\mu}\langle\Delta s,D_{ss}^{2}\eta_{\rho}(\hat{x},s;\mu)\Delta s\rangle},$ (77b)
$\delta_{\rho}(\hat{x},s;\mu)$ $=\sqrt{(\delta_{\hat{x},\rho}(\hat{x},s;\mu))^{2}+(\delta_{s,\rho}(\hat{x},s;\mu))^{2}}=\|\Delta w\|_{\eta_{\rho},w},$ (77c)
$\xi_{\hat{x},\rho}(\hat{x},s;\mu)$ $=\sqrt{\frac{1}{\mu}\langle\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu),D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)^{-1}\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)\rangle},$ (77d)
$\xi_{s,\rho}(\hat{x},s;\mu)$ $=\sqrt{-\frac{1}{\mu}\langle\nabla_{s}\eta_{\rho}(\hat{x},s;\mu),D^{2}_{ss}\eta_{\rho}(\hat{x},s;\mu)^{-1}\nabla_{s}\eta_{\rho}(\hat{x},s;\mu)\rangle},$ (77e)
$\xi_{\rho}(\hat{x},s;\mu)$ $=\sqrt{(\xi_{\hat{x},\rho}(\hat{x},s;\mu))^{2}+(\xi_{s,\rho}(\hat{x},s;\mu))^{2}}=\|\nabla_{w}\eta_{\rho}(\hat{x},s;\mu)\|^{*}_{\eta_{\rho},w}.$ (77f)
These quantities serve as surrogate measures of the quality of the current iterate.
Remark 3
Note that $\Delta\hat{x}=\mathcal{B}^{*}\Delta x$, $\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)=\mathcal{B}^{*}\nabla_{x}\eta_{\rho}(\hat{x},s;\mu)$, and
$D_{xx}^{2}\eta_{\rho}(\hat{x},s;\mu)=\mathcal{B}\,D_{\hat{x}\hat{x}}^{2}\eta_{\rho}(\hat{x},s;\mu)\,\mathcal{B}^{*}.$
This implies
$\delta_{x,\rho}(x,s;\mu)$ $=\sqrt{\frac{1}{\mu}\langle\Delta x,{D_{xx}^{2}\eta_{\rho}(\hat{x},s;\mu)}\Delta x\rangle}=\delta_{\hat{x},\rho}(\hat{x},s;\mu),$ (78)
$\xi_{x,\rho}(x,s;\mu)$ $=\sqrt{\frac{1}{\mu}\langle\nabla_{x}\eta_{\rho}(\hat{x},s;\mu)D^{2}_{xx}\eta_{\rho}(\hat{x},s;\mu)^{-1}\nabla_{x}\eta_{\rho}(\hat{x},s;\mu)\rangle}$
$\ =\xi_{\hat{x},\rho}(\hat{x},s;\mu)$
In practical computations, it is unnecessary to explicitly form $\Delta\hat{x}$ and $\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)$. The quantities $\delta_{x,\rho}(x,s;\mu)$ and $\xi_{x,\rho}(x,s;\mu)$ can be computed directly from $\Delta x$ and $\nabla_{x}\eta_{\rho}(\hat{x},s;\mu)$. For the theoretical analysis, however, we work with $\delta_{\hat{x},\rho}(\hat{x},s;\mu)$ and $\xi_{\hat{x},\rho}(\hat{x},s;\mu)$.
Define the following auxiliary merit functions:
$\tilde{\delta}_{\hat{x},\rho}(\hat{x},s;\mu)$ $=\sqrt{\dfrac{1}{\mu}\langle\widetilde{\Delta\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)\widetilde{\Delta\hat{x}}\rangle}$, (79a)
$\tilde{\delta}_{s,\rho}(\hat{x},s;\mu)$ $=\sqrt{-\dfrac{1}{\mu}\langle\widetilde{\Delta s},D^{2}_{ss}\eta_{\rho}(\hat{x},s;\mu)\widetilde{\Delta s}\rangle}$, (79b)
$\tilde{\delta_{\rho}}(\hat{x},s;\mu)$ $=\sqrt{(\tilde{\delta}_{\hat{x},\rho}(\hat{x},s;\mu))^{2}+(\tilde{\delta}_{s,\rho}(\hat{x},s;\mu))^{2}}=\|\widetilde{\Delta w}\|_{\eta_{\rho},w},$ (79c)
[…]
If $\xi_{\rho}(\hat{x},s;\mu)=0$, then $(\hat{x},s)$ is the saddle point of the minimax problem (33). This defines a central path that coincides with the one generated by IPMs, as shown in Theorem $3.$ Specifically,
$\mathcal{A}x=b,\,\mathcal{A}^{*}\lambda+s=c,\,x\in\operatorname{int}\,(\mathbb{K}),\,s\in\operatorname{int}\,(\mathbb{K}),\,x\circ s=\mu e$ (80)
$\Longleftrightarrow$ $\quad\mathcal{A}x=b,\,\mathcal{A}^{*}\lambda+s=c,\,\xi_{\rho}(\hat{x},s;\mu)=0.$
Motivated by this, we define the central-path neighborhood for the SNM based on the reduced SBAL function $\eta_{\rho}$ by
$\mathcal{N}(\kappa,\mu,\rho)=\left\{(x,s,\lambda)\in\mathbb{E}\times\mathbb{E}\times\mathbb{R}^{m}\,|\,\mathcal{A}x=b,\,\mathcal{A}^{*}\lambda+s=c,\,\xi_{\rho}(\hat{x},s;\mu)\leq\kappa\right\},$ (81)
where $\kappa=0.1$ is fixed throughout the algorithm and the complexity analysis.
This neighborhood differs from the standard neighborhoods used in classical interior-point Monteiro and Zhang (1998); Nesterov and Todd (1998); Schmieta and Alizadeh (2003) or non-interior Burke and Xu (2000, 1998); Chen and Tseng (2003); Zhao and Li (2003) path-following methods in the literature. It is defined via the merit function induced by the minimax problem, and provides a more faithful characterization of the behavior of the SNM. The proposed method follows the standard paradigm of path-following methods. In the first phase, the iterates are driven into $\mathcal{N}(\kappa,\mu^{(0)},\rho)$. In the second phase, the trajectory is confined to $\mathcal{N}(\kappa,\mu^{(k)},\rho)$, which ensures that all subsequent iterates remain well behaved.
4.2 Two-phase path-following framework
Before introducing the two-phase framework, we present some necessary notations. Let $v:=(x,s,\lambda)$. Define $\Delta v:=(\Delta x,\Delta s,\Delta\lambda)$ as the concatenation of $\Delta x$, $\Delta s$, and $\Delta\lambda$. Let $K(\theta_{\rho}):=\bigl\{(\hat{x},s):\theta_{\rho}(\hat{x},s;\mu)<+\infty\bigr\}$. For a given initialization point $w^{(0,0)}:=(\hat{x}^{(0,0)},s^{(0,0)})\in K(\theta_{\rho})$, let $\bar{x}\in\mathbb{E}$ with $\mathcal{A}\bar{x}=b$, and $x^{(0,0)}=\bar{x}+\mathcal{B}\hat{x}^{(0,0)}$. For $t\in[0,1]$, define the perturbed function
$\eta_{t,\rho}(w;\mu^{(0)}):=\eta_{\rho}(w;\mu^{(0)})-t\langle\nabla_{w}\eta_{\rho}(w^{(0,0)};\mu^{(0)}),w\rangle.$ (82)
Since $\eta_{\rho}$ is a nondegenerate $\mu$-self-concordant convex-concave function, $\eta_{t,\rho}$ inherits the same property for any $t\geq 0$.
The first-phase algorithm aims to generate a strictly feasible point within the neighborhood $\mathcal{N}(\kappa,\mu^{(0)},\rho)$. The parameter $t$ is used to scale the merit function so that the initial point $(x^{(0,0)},s^{(0,0)},\lambda^{(0,0)})$ lies in the central-path neighborhood. For this purpose, consider the minimax problem
$\min_{\hat{x}}\max_{s}\left\{\eta_{t,\rho}(w;\mu^{(0)})\right\},$
with first-order optimality conditions
$\mathcal{B}^{*}(c-y_{\rho})-t\mathcal{B}^{*}(c-y_{\rho}^{(0,0)})=0,$ (83)
$z_{\rho}-x-t(z_{\rho}^{(0,0)}-x^{(0,0)})=0,$
where $y_{\rho}^{(0,0)}=y_{\rho}(x^{(0,0)},s^{(0,0)};\mu^{(0)})$ and $z_{\rho}^{(0,0)}=z_{\rho}(x^{(0,0)},s^{(0,0)};\mu^{(0)})$.
Applying Newton’s method to the nonlinear system (83) leads to the search direction $(\Delta\hat{x},\Delta s)$ satisfying
$\begin{pmatrix}\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}&-\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\\ -\mathcal{H}^{-1}\mathcal{W}\mathcal{B}&-\rho^{-1}\mathcal{H}^{-1}\end{pmatrix}\begin{pmatrix}\Delta\hat{x}\\ \Delta s\end{pmatrix}=-\begin{pmatrix}\mathcal{B}^{*}(c-y_{\rho})\\ z_{\rho}-x\end{pmatrix}+t\begin{pmatrix}\mathcal{B}^{*}(c-y_{\rho}^{(0,0)})\\ z_{\rho}^{(0,0)}-x^{(0,0)}\end{pmatrix}$. (84)
In practical computations, the variable $\hat{x}$ is neither explicitly computed nor formulated. Following Theorem 3.4, we instead solve the system in $(x,s,\lambda)$:
$\begin{pmatrix}\mathcal{A}&0&0\\ 0&\mathcal{I}_{\mathbb{E}}&\mathcal{A}^{*}\\ \mathcal{H}^{-1}\mathcal{W}&\rho^{-1}\mathcal{H}^{-1}&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta s\\ \Delta\lambda\end{pmatrix}=-\begin{pmatrix}\mathcal{A}x-b\\ \mathcal{A}^{*}\lambda+s-c\\ x-z_{\rho}\end{pmatrix}+t\begin{pmatrix}0\\ \mathcal{A}^{*}\lambda^{(0,0)}+s^{(0,0)}-c\\ x^{(0,0)}-z_{\rho}^{(0,0)}\end{pmatrix}$. (85)
The resulting search direction coincides with that obtained by solving the reduced system (84) under the constraint $\mathcal{A}x=b$. This equivalence is formalized in the following corollary.
Corollary 2
Let $(x,s,\lambda)$ satisfy the linear constraint $\mathcal{A}x=b$. For the given point $(x^{(0,0)},s^{(0,0)},\lambda^{(0,0)})$ and any $\mu^{(0)}>0,\rho\geq 1$, suppose that $(\Delta x,\Delta s,\Delta\lambda)$ solves (85). Then the pair $(\Delta\hat{x},\Delta s):=(\mathcal{B}^{*}\Delta x,\Delta s)$ is the unique solution of (84). Conversely, if $(\Delta\hat{x},\Delta s)$ solves (84), define
$\Delta\lambda:=\left({\mathcal{A}\mathcal{A}^{*}}\right)^{-1}\mathcal{A}(-s-\Delta s+c-\mathcal{A}^{*}\lambda+t(\mathcal{A}^{*}\lambda^{(0,0)}+s^{(0,0)}-c)).$ (86)
Then $(\Delta x,\Delta s,\Delta\lambda):=(\mathcal{B}\Delta\hat{x},\Delta s,\Delta\lambda)$ is the unique solution of (85).
Proof
The proof follows the same arguments as in Theorem 3.4, with a minor modification due to the shift term. Details are omitted.
We start from the initial point $(x^{(0,0)},s^{(0,0)},\lambda^{(0,0)})$ satisfying $\mathcal{A}\bar{x}=b$. Choose $t^{(0)}$ such that
$(1-t^{(0)})\delta_{\rho}(w^{(0,0)};\mu^{(0)})\leq\frac{\kappa}{2}.$ (87)
Repeatedly solving the linear system (85) while updating $t$, we eventually obtain a point $(x^{(0,j)},s^{(0,j)},\lambda^{(0,j)})$ lying in the neighborhood $\mathcal{N}(\kappa,\mu^{(0)},\rho)$. The corresponding algorithmic framework is summarized below.
Algorithm 1 The first phase of PFSNM
Step 1: Choose $(\hat{x}^{(0,0)},s^{(0,0)})\in K(\theta_{\rho})$, $\lambda^{(0,0)}\in\mathbb{R}^{m}$, $\rho\geq 1,\mu^{(0)}>0$, and $t^{(0)}\in[0,1]$
     satisfying (87). Set $j:=0$.
Step 2: Compute $\delta_{\rho}^{(0,j)}:=\delta_{\rho}({w}^{(0,j)};\mu^{(0)})$. If $\delta^{(0,j)}_{\rho}\leq\kappa$, compute the Newton
     direction $\Delta v^{(0,j)}$ from (45), set
$v^{(0)}=v^{(0,j)}+\Delta v^{(0,j)},$
     and terminate. Otherwise go to Step 3.
Step 3: Update $t^{(j+1)}=(1-\alpha^{(j)})t^{(j)}$, where
$\alpha^{(j)}=\min\left\{\frac{\kappa}{4t^{(j)}{\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)};\mu^{(0)})^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)};\mu^{(0)})}\|_{\eta_{\rho},w^{(0,j)}}},1\right\}.$ (88)
Step 4: Compute the search direction $\Delta v^{(j)}$ from (85) with $t=t^{(j+1)}$. Set
$v^{(0,j+1)}=v^{(0,j)}+\Delta v^{(j)}.$ (89)
     Set $j:=j+1$ and return to Step 2.
Throughout Algorithm 1, the iterates $(x^{(0,j)},s^{(0,j)},\lambda^{(0,j)})$ always satisfy the primal constraint by (85). When Algorithm 1 terminates, the second equation of (45) yields
$\mathcal{A}^{*}\lambda^{(0)}+s^{(0)}-c=0.$
Moreover, Theorem 2.3(iii) gives
$\xi_{\rho}(w^{(0)};\mu^{(0)})\leq\dfrac{\delta^{(0,j)}_{\rho}}{2}\leq\kappa,$ (90)
which implies that $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$.
Once a point in $\mathcal{N}(\kappa,\mu^{(0)},\rho)$ has been obtained, we decrease the barrier parameter and, for each new value of the parameter, apply Newton steps to re-enter the corresponding neighborhood. The Newton direction is generated by (45). In what follows, we present the second-phase algorithm.
Algorithm 2 The second phase of PFSNM
Step 1: Input $v^{(0,0)}:=v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$, $\rho\geq 1$, and $\mu^{(0)}>0$. Set $k:=0$ and $j:=0$.
Step 2: Compute the search direction $\Delta v^{(k,j)}$ from (45) and update
$v^{(k,j+1)}=v^{(k,j)}+\Delta v^{(k,j)}.$
Step 3: Compute $\xi^{(k,j+1)}_{\rho}:=\xi_{\rho}(\hat{x}^{(k,j+1)},s^{(k,j+1)};\mu^{(k)})$. If $\xi^{(k,j+1)}_{\rho}>\kappa$, set $j:=j+1$
     and go to Step 2. Otherwise, go to Step 4.
Step 4: Set $v^{(k)}:=v^{(k,j+1)}$ and $\xi^{(k)}_{\rho}:=\xi^{(k,j+1)}_{\rho}$. If $\mu^{(k)}>\varepsilon$, update $\mu^{(k+1)}=\sigma\mu^{(k)}$,
     $v^{(k+1,0)}=v^{(k)}$, set $k:=k+1$, $j:=0$, and return to Step 2. Otherwise, terminate
     and return $v^{(k)}$.
After each reduction of the barrier parameter, the current iterate $v^{(k)}$ is used as the starting point for the next outer iteration and is recentered until it re-enters the new neighborhood $\mathcal{N}(\kappa,\mu^{(k+1)},\rho)$. The well-definedness of this procedure and the bound on the number of inner Newton steps will be established in the next section.
Some remarks on Algorithms 1 and 2 are in order:
  1. (i)
    In practical computations, the variable $\hat{x}$ is never explicitly formed. Computations are performed directly with the primal variable $x$ since all the quantities can be evaluated from $(x,s,\lambda)$; see Remark 3 for details.
  2. (ii)
    The iterate entering Algorithm 2 satisfies the primal and dual constraints. Consequently, by (45), (85), and the iteration scheme, all subsequent iterates preserve feasibility:
    $\mathcal{A}x^{(k,j)}$ $=b,\,$ $\mathcal{A}^{*}\lambda^{(k,j)}+s^{(k,j)}=c,\quad\forall\,j\geq 0,\,k\geq 1.$ (91)
  3. (iii)
    Once Algorithm 2 terminates, we have
    $\mu^{(k)}\leq\varepsilon\;\text{ and }\;\xi^{(k)}_{\rho}=\|\nabla_{w}\eta_{\rho}(w^{(k)};\mu^{(k)})\|_{\eta_{\rho},w^{(k)},\mu^{(k)}}^{*}\leq\kappa.$ (92)
    It follows from the definition of $\xi^{(k)}_{\rho}$ and Corollary 1 that
    $\|z^{(k)}_{\rho}-x^{(k)}\|$ $\leq\sqrt{\langle z^{(k)}_{\rho}-x^{(k)},\mathcal{H}(z^{(k)}_{\rho}-x^{(k)})\rangle}$ (93)
    $\leq\sqrt{\frac{\mu^{(k)}}{\rho}}\|\nabla_{w}\eta_{\rho}(w^{(k)},\mu^{(k)})\|^{*}_{\eta_{\rho},w^{(k)},\mu^{(k)}}$
    $\leq\sqrt{\frac{\varepsilon}{\rho}}\kappa.$
From remarks (ii) and (iii), when Algorithm 2 terminates, an approximate KKT solution $(x^{(k)},s^{(k)},\lambda^{(k)})$ to the original SCP problem (1) is obtained, satisfying
$\mathcal{A}x^{(k)}=b,\;\mathcal{A}^{*}\lambda^{(k)}+s^{(k)}=c,\;\|\Phi_{\rho}(x^{(k)},s^{(k)};\mu^{(k)})\|\leq 2\sqrt{\frac{\varepsilon}{\rho}}\kappa.$ (94)
4.3 Linear system in the algorithm
During practical computations, the primal and dual constraints are satisfied at all iterates (see (91)). Thus, the Newton systems (85) and (45) arising in Algorithms 1 and 2 can be written in the following unified form:
$\begin{pmatrix}\mathcal{A}&0&0\\ 0&\mathcal{I}_{\mathbb{E}}&\mathcal{A}^{*}\\ \mathcal{H}^{-1}\mathcal{W}&\rho^{-1}\mathcal{H}^{-1}&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta s\\ \Delta\lambda\end{pmatrix}=\begin{pmatrix}0\\ 0\\ r\end{pmatrix}$, (95)
where
[…]
Directly solving the full Newton system (95) is computationally expensive, particularly for large-scale problems where the dimensions of primal and dual variables $x$ and $s$ are significantly larger than the number of constraints. To reduce the computational cost, we eliminate $\Delta x$ and $\Delta s$ from (95), which leads to the following reduced system:
$\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}\Delta\lambda$ $=-\rho\mathcal{A}\mathcal{W}^{-1}\mathcal{H}r,$ (96a)
$\Delta s$ $=-\mathcal{A}^{*}\Delta\lambda,$ (96b)
$\Delta x$ $=\mathcal{W}^{-1}(\mathcal{H}r-\rho^{-1}\Delta s).$ (96c)
For both IPMs and classical SNMs, the dominant computational cost typically comes from forming and factorizing the Schur complement $\mathcal{A}\mathcal{D}\mathcal{A}^{*}$. In our method, the matrix takes the form $\mathcal{D}=\mathcal{W}^{-1}$, where $\mathcal{W}$ is an iterate-dependent matrix determined by the barrier function in PFSNM. Accordingly, improving the efficiency of constructing $\mathcal{A}\mathcal{D}\mathcal{A}^{*}$ is crucial for large-scale performance. To this end, Proposition 3 provides closed-form expressions for $\mathcal{D}$ (and hence for $\mathcal{A}\mathcal{D}\mathcal{A}^{*}$) in PFSNM for the three most common symmetric cones.
Proposition 3
Let $\mathbb{K}$ be one of the symmetric cones listed below, and let $e$ denote the corresponding Jordan identity element. Then the corresponding Schur complement admits the following explicit representations.
  1. (i)
    Let $\mathbb{K}=\mathbb{R}^{n}_{+}$. Then, for any $z\in\mathbb{R}^{n}_{++}$, $\phi(z)=-\sum_{i=1}^{n}\ln\,z_{i}$, and
    $\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}=\frac{\rho}{\mu}\mathcal{A}\big(\operatorname{Diag}(z_{\rho})\big)^{2}\mathcal{A}^{*}.$
  2. (ii)
    Let $\mathbb{K}=\mathbb{Q}^{n+1}$. Then, for any $z\in\operatorname{int}\,(\mathbb{Q}^{n+1})$, $\phi(z)=-\frac{1}{2}\ln\big(z_{0}^{2}-\|\bar{z}\|^{2}\big),$ and
    $\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}=\frac{\rho}{\mu}\Big(\det(z_{\rho})\,\mathcal{A}\mathcal{A}^{*}+2(\mathcal{A}z_{\rho})(\mathcal{A}z_{\rho})^{*}-2\det(z_{\rho})\,(\mathcal{A}e)(\mathcal{A}e)^{*}\Big).$
  3. (iii)
    Let $\mathbb{K}=\mathbb{S}^{n}_{+}$. Then, for any $Z\in\mathbb{S}^{n}_{++}$, $\phi(Z)=-\ln\det(Z)$, and
    $\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}=\frac{\rho}{\mu}\,\mathcal{A}(Z_{\rho}\otimes_{s}Z_{\rho})\mathcal{A}^{*}.$
Proof
By definition, ${\mathcal{W}}=\dfrac{\mu}{\rho}D^{2}\phi(z_{\rho})$. The results follow directly from the explicit formulas (see (Vieira, 2007, Proposition 2.6.1)) of $D^{2}\phi$.
The importance of having explicit formulas has already been observed in the SDP literature. The SNM in Chen and Tseng (2003), based on the smoothing FB function, has a Schur-complement formation cost comparable to that of the most expensive AHO direction in IPMs. If the smoothing CHKS function is used instead, the formation cost becomes cheaper than that of AHO, but remains more expensive than that of the NT direction. In our earlier work Zhang et al. (2024), we showed that once explicit formulas of the Schur complement are available, the formation cost can be reduced to the same order as that of the NT direction. Compared with Zhang et al. (2024), the work in this paper leverages self-concordant properties to simplify the derivation of this explicit Schur complement formation, resulting in a more direct construction.
For SOCP, the advantage of explicit Schur complement formation is even more significant. As shown in case (ii) of Proposition 3, the Schur complement takes the form of a scaled matrix combined with two rank-one updates. However, the vector $u:=\mathcal{A}z_{\rho}$ appearing in these rank-one terms is typically dense, which causes the full Schur complement to be dense as well. Figure 2 illustrates this structure and shows how these components combine to yield a fully dense matrix. This density poses a major computational challenge in large-scale settings. Even accelerating the evaluation of $\mathcal{D}$ as described in Fukushima et al. (2002) does not solve the problem, because the Schur complement remains dense in any case. The explicit representation in Proposition 3 avoids forming this dense matrix entirely. The terms $\mathcal{A}\mathcal{A}^{*}$ and $\mathcal{A}e$ depend only on the problem data and can be precomputed once. Each subsequent iteration then requires only one matrix-vector product $u=\mathcal{A}z_{\rho}$, along with simple low-rank updates and the scalar $\det(z_{\rho})$. With this structure, one can apply the product-form Cholesky factorization approach in Alizadeh and Goldfarb (2003) or the expanded sparse representation technique in Zhang et al. (2026). Both approaches exploit the low-rank structure and avoid forming a dense Schur complement.
[…]
Figure 2: Structure of the SOCP Schur complement in PFSNM
5 Complexity analysis
In this section, we establish a polynomial iteration-complexity bound for PFSNM for SCP. The analysis is divided into two parts. We first derive a complexity bound for the first phase. Starting from the point produced by this phase, Algorithm 2 then attains a complexity bound of order $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, matching the classical short-step interior-point complexity Vavasis and Ye (1996).
The complexity bound for the first phase is presented in the following theorem.
Theorem 5.1
Suppose that $(\hat{x}^{(0,0)},s^{(0,0)})\in K(\theta_{\rho})$ and that $t^{(0)}$ in Algorithm 1 satisfies (87). Then Algorithm 1 requires at most
$\mathcal{O}\left(\ln\left(1+\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}\right)\right)$ (97)
iterations to attain a starting point $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$, where $\Theta_{1}(\cdot)$ is a properly chosen universal positive continuous and nondecreasing function on $\mathbb{R}_{+}$.
Proof
Since $\mu$ remains fixed in the first phase, write $\eta_{\rho}(w)$ and $\eta_{t,\rho}(w)$ in place of $\eta_{\rho}(w;\mu^{(0)})$ and $\eta_{t,\rho}(w;\mu^{(0)})$, respectively. Define the merit function associated with $\eta_{t,\rho}$ by
$\delta_{t,\rho}(w):=\|(D^{2}_{ww}\eta_{\rho}(w))^{-1}\nabla_{w}\eta_{t,\rho}(w)\|_{\eta_{t,\rho},w}=\|(D^{2}_{ww}\eta_{\rho}(w))^{-1}\nabla_{w}\eta_{t,\rho}(w)\|_{\eta_{\rho},w}.$ (98)
Noting that $\nabla_{w}\eta_{t,\rho}(w)=\nabla_{w}\eta_{\rho}(w)-t\nabla_{w}\eta_{\rho}(w^{(0,0)})$,
$\delta_{t^{(0)},\rho}(w^{(0,0)})=(1-t^{(0)})\delta_{\rho}(w^{(0,0)}).$ (99)
We now prove by induction that $\delta_{t^{(j)},\rho}(w^{(0,j)})\leq\frac{\kappa}{2}$ at each iteration $j$. For $j=0$, the claim follows from (87) and (99). Assume that $\delta_{t^{(j)},\rho}(w^{(0,j)})\leq\frac{\kappa}{2}$ at the $j$-th iteration. Then
$\delta_{t^{(j+1)},\rho}(w^{(0,j)})-\delta_{t^{(j)},\rho}(w^{(0,j)})$ (100)
$=$ $\ \|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{t^{(j+1)},\rho}(w^{(0,j)})\|_{\eta_{\rho,w}}$
$\qquad\qquad\qquad-\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{t^{(j)},\rho}(w^{(0,j)})\|_{\eta_{\rho},w}$
$\overset{(i)}{\leq}$ $\ \alpha^{(j)}t^{(j)}\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)})\|_{\eta_{\rho},w^{(0,j)}}$
$\overset{(ii)}{\leq}$ $\ \frac{\kappa}{4}.$
Here, $(i)$ follows from the identity for $\nabla_{w}\eta_{t,\rho}(w)$ and that $\|\cdot\|_{\eta_{\rho},w^{(0,j)}}$ defines a norm. Inequality $(ii)$ follows from (88). Thus,
$\delta_{t^{(j+1)},\rho}(w^{(0,j)})\leq\frac{\kappa}{2}+\frac{\kappa}{4}=\frac{3\kappa}{4}<\kappa.$ (101)
By (101) and Theorem 2.3 (ii)–(iii), we have
$\delta_{t^{(j+1)},\rho}(w^{(0,j+1)})\leq\left(\frac{\delta_{t^{(j+1)},\rho}(w^{(0,j)})}{1-\delta_{t^{(j+1)},\rho}(w^{(0,j)})}\right)^{2}\leq\frac{\kappa}{2},$ (102)
which completes the induction argument. Hence,
$\delta_{t^{(j)},\rho}(w^{(0,j)})\leq\frac{\kappa}{2},\quad\forall\,j\geq 0.$ (103)
It follows from (Nemirovski, 1999, Lemma 8.3(a)), there exists a universal positive continuous and nondecreasing function $\Theta_{1}$ such that
$\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)})\|_{\eta_{\rho},w^{(0,j)}}\leq\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)})).$ (104)
Consequently, by (88),
$\underline{\alpha}:=\min\left\{1,\frac{\kappa}{4t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}\right\}\leq\alpha^{(j)}\leq 1.$ (105)
We now consider two different cases associated with the value of $\underline{\alpha}$.
If $\underline{\alpha}=1$, then $\alpha^{(0)}=1$ and $t^{(1)}=0$. Since $\delta_{t^{(j)},\rho}(w^{(0,j)})\leq\frac{\kappa}{2}$ at each iteration $j$, we have
$\delta_{\rho}(w^{(0,1)})=\delta_{t^{(1)},\rho}(w^{(0,1)})\leq\frac{\kappa}{2}.$ (106)
Together with (90), this implies that one additional iteration leads to $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$. Suppose that $\underline{\alpha}<1$. Then,
$t^{(j)}\leq(1-\underline{\alpha})^{j}t^{(0)}.$ (107)
Combining (104) and (107) gives
$\delta_{\rho}(w^{(0,j)};\mu^{(0)})$ (108)
$=$ $\ \|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,j)})\|_{\eta_{\rho},w^{(0,j)}}$
$\leq$ $\ \delta_{t^{(j)},\rho}(w^{(0,j)})+t^{(j)}\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)})\|_{\eta_{\rho},w^{(0,j)}}$
$\leq$ $\ \frac{\kappa}{2}+(1-\underline{\alpha})^{j}t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)})).$
Using (90), it requires at most $\mathcal{O}\left(\ln\left(1+\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}\right)\right)$ iterations to attain the starting point $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$. This completes the proof.
Having established the bound for the first-phase of PFSNM, we now turn to analyze the complexity bound of Algorithm 2. As a preliminary step, Theorems 5.2 and 5.3 quantify the effect of updates in the barrier parameter $\mu$ on both $\nabla_{w}\eta_{\rho}(\hat{x},s;\mu)$ and $S_{\eta_{\rho}}(\hat{x},s;\mu)$. For notational simplicity, we omit the arguments $(\hat{x},s;\mu)$ and write, for example, $\nabla_{w}\eta_{\rho}:=\nabla_{w}\eta_{\rho}(\hat{x},s;\mu)$ whenever first- or second-order derivatives of $\eta_{\rho}$ with respect to $w=(\hat{x},s)$ are mentioned. All derivatives with respect to $\mu$ are denoted by a prime.
Theorem 5.2
For any $\mu>0$, $\rho\geq 1$, any direction $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, and any point $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\left|\langle h,\nabla_{w}\eta_{\rho}^{\prime}(\hat{x},s;\mu)\rangle\right|\leq\sqrt{\dfrac{2\nu}{\mu}}\sqrt{S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]}.$ (109)
Proof
Combining Theorem 3.1 and Corollary 1 yields
$\nabla_{w}\eta_{\rho}^{\prime}(\hat{x},s;\mu)=\begin{pmatrix}-\mathcal{B}^{*}y^{\prime}_{\rho}\\ z^{\prime}_{\rho}\end{pmatrix}=\begin{pmatrix}\mathcal{B}^{*}\mathcal{H}^{-1}\nabla\phi(z_{\rho})\\ -\rho^{-1}\mathcal{H}^{-1}\nabla\phi(z_{\rho})\end{pmatrix}$. (110)
For any $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\langle\nabla_{w}\eta_{\rho}^{\prime}(\hat{x},s;\mu),h\rangle=\langle\mathcal{B}h_{\hat{x}},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle-\rho^{-1}\langle h_{s},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle.$ (111)
Let $h_{x}=\mathcal{B}h_{\hat{x}}$. Then,
$|\langle\nabla_{w}\eta^{\prime}_{\rho},h\rangle|$ $\leq\left|\langle h_{x},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle\right|+\rho^{-1}\left|\langle h_{s},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle\right|$ (112)
$=\left|\langle h_{x},\mathcal{H}^{-\frac{1}{2}}\mathcal{W}^{\frac{1}{2}}\mathcal{W}^{-\frac{1}{2}}\mathcal{H}^{-\frac{1}{2}}\nabla\phi(z_{\rho})\rangle\right|+\rho^{-1}\left|\langle h_{s},\mathcal{H}^{-\frac{1}{2}}\mathcal{H}^{-\frac{1}{2}}\nabla\phi(z_{\rho})\rangle\right|$
$\leq\sqrt{\langle h_{x},\mathcal{H}^{-1}\mathcal{W}h_{x}\rangle}\sqrt{\langle\nabla\phi(z_{\rho}),\mathcal{W}^{-1}\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle}$
$\qquad\qquad\qquad+\rho^{-1}\sqrt{\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle}\sqrt{\langle\nabla\phi(z_{\rho}),\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle}$
$\overset{(i)}{\leq}\sqrt{\dfrac{\langle h_{\hat{x}},D_{\hat{x}\hat{x}}^{2}\eta_{\rho}h_{\hat{x}}\rangle}{\rho}}\sqrt{\dfrac{\rho}{\mu}\langle\nabla\phi(z_{\rho}),(D^{2}\phi(z_{\rho}))^{-1}\nabla\phi(z_{\rho})\rangle}$
$\qquad\qquad\qquad+\sqrt{-\dfrac{\langle h_{s},D_{ss}^{2}\eta_{\rho}h_{s}\rangle}{\rho}}\sqrt{\dfrac{\rho}{\mu}\langle\nabla\phi(z_{\rho}),(D^{2}\phi(z_{\rho}))^{-1}\nabla\phi(z_{\rho})\rangle}$
$\overset{(ii)}{=}\sqrt{\dfrac{\nu}{\mu}}\left(\sqrt{\langle h_{\hat{x}},D_{\hat{x}\hat{x}}^{2}\eta_{\rho}h_{\hat{x}}\rangle}+\sqrt{-\langle h_{s},D_{ss}^{2}\eta_{\rho}h_{s}\rangle}\right)$
$\leq\sqrt{\dfrac{2\nu}{\mu}}\sqrt{S_{\eta_{\rho}}[h,h]}.$
Here, $(i)$ follows from Corollary 1, $\mathcal{H}^{-1}\prec\mathcal{I}_{\mathbb{E}}$, and $\mathcal{H}^{-1}\prec\mathcal{W}^{-1}$. The identity $(ii)$ follows from (11).
Theorem 5.3
For any $\mu>0$, $\rho\geq 1$, any direction $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, and any point $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\left|S_{\eta_{\rho}}^{\prime}(\hat{x},s;\mu)[h,h]\right|\leq\frac{\rho(1+{2\sqrt{\nu}})}{\mu}S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h].$ (113)
Proof
Let $h_{x}=\mathcal{B}h_{\hat{x}}$ and $\hat{h}=(\hat{h}_{\hat{x}},\hat{h}_{s})=(\mathcal{H}^{-1}h_{x},\mathcal{H}^{-1}h_{s})$. Define
$\omega(\mu):=S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h].$
By Corollary 1,
$\omega(\mu)=\omega_{x}(\mu)+\omega_{s}(\mu),$
where
$\omega_{x}(\mu):=\rho\langle h_{x},(\mathcal{I}_{\mathbb{E}}-\mathcal{H}^{-1})h_{x}\rangle,\quad\omega_{s}(\mu):=\rho^{-1}\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle.$
Differentiating $\omega_{x}(\mu)$ with respect to $\mu$ yields
$\omega^{\prime}_{x}(\mu)$ $=\langle\mathcal{H}^{-1}h_{x},D^{2}\phi(z_{\rho})\mathcal{H}^{-1}h_{x}\rangle+\mu D^{3}\phi(z_{\rho})[z_{\rho}^{\prime},\mathcal{H}^{-1}h_{x},\mathcal{H}^{-1}h_{x}]$ (114)
$=\frac{\rho}{\mu}\langle h_{x},\mathcal{H}^{-1}\mathcal{W}\mathcal{H}^{-1}h_{x}\rangle+\mu D^{3}\phi(z_{\rho})[z_{\rho}^{\prime},\hat{h}_{\hat{x}},\hat{h}_{\hat{x}}$]
$\overset{(i)}{\leq}\frac{\rho}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]+2\mu\sqrt{D^{2}\phi(z_{\rho})[z_{\rho}^{\prime},z_{\rho}^{\prime}]}D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}},\hat{h}_{\hat{x}}$]
$=\frac{\rho}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]+2\sqrt{D^{2}\phi(z_{\rho})[z_{\rho}^{\prime},z_{\rho}^{\prime}]}\left<\mathcal{H}^{-1}h_{x},\mathcal{W}\mathcal{H}^{-1}h_{x}\right$>
$\overset{(ii)}{\leq}\frac{\rho}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]+2\sqrt{\frac{\rho}{\mu}\langle\nabla\phi(z_{\rho}),\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}$]
$\overset{(iii)}{\leq}\frac{\rho}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]+\frac{2\rho}{\mu}\sqrt{\langle\nabla\phi(z_{\rho}),(D^{2}\phi(z_{\rho}))^{-1}\nabla\phi(z_{\rho})\rangle}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}$]
$\overset{(iv)}{=}\dfrac{\rho(1+2\sqrt{\nu})}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}$].
Here, the inequalities (i)(i)$(ii)$ follow from Corollary 1, Theorem 3.1, the 11-self-concordance of $\phi$, and $\mathcal{H}^{-1}\prec\mathcal{I}_{\mathbb{E}}$. The inequality $(iii)$ uses $\mathcal{H}^{-1}\prec\mathcal{W}^{-1}$. The final identity $(iv)$ is due to (11).
Similarly, we obtain
$\omega^{\prime}_{s}(\mu)\leq-\frac{1+2\sqrt{\nu}}{\mu}D^{2}_{ss}\eta_{\rho}[h_{s},h_{s}].$ (115)
Combing (114) and (115) gives
$\omega^{\prime}(\mu)$ $\leq\dfrac{\rho(1+2\sqrt{\nu})}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]-\frac{1+2\sqrt{\nu}}{\mu}D^{2}_{ss}\eta_{\rho}[h_{s},h_{s}]$ (116)
$\leq\frac{\rho(1+{2\sqrt{\nu}})}{\mu}S_{\eta_{\rho}}[h,h],$
which completes the proof.
The following theorem provides an upper bound for the primal-dual gap function $\theta_{\rho}(\hat{x},s;\mu)$ that depends only on $\kappa$ and $\mu$.
Theorem 5.4
For any $\mu>0$, $\rho\geq 1$, and any point $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$, suppose that
$\xi_{\rho}(\hat{x},s;\mu)\leq\kappa.$
Then the primal-dual gap function satisfies
$\theta_{\rho}(\hat{x},s;\mu)\leq\kappa\mu.$ (117)
Proof
To quantify the gaps between the current value $\eta_{\rho}$ and the primal and dual optimal values, define
$\theta_{\hat{x},\rho}(\mu)$ $=\eta_{\rho}(\hat{x},{s};\mu)-\eta_{\rho}(\hat{x}_{\rho}(s,\mu),{s};\mu),$ (118)
$\theta_{s,\rho}(\mu)$ $=\eta_{\rho}(\hat{x},{s}_{\rho}(\hat{x},\mu);\mu)-\eta_{\rho}(\hat{x},{s};\mu).$
If $\xi_{\rho}(\hat{x},s;\mu)\leq\kappa$, then by Theorem 2.3(v),
$\max\left\{\tilde{\delta}_{\hat{x},\rho}(\hat{x},s;\mu),\tilde{\delta}_{s,\rho}(\hat{x},s;\mu)\right\}\leq 2\kappa.$ (119)
Consequently, we have
$\theta_{s,\rho}(\mu)$ $=\eta_{\rho}(\hat{x},{s}(\hat{x},\mu);\mu)-\eta_{\rho}(\hat{x},{s};\mu)$ (120)
$=-\int_{0}^{1}\langle\widetilde{\Delta s},\nabla_{s}\eta_{\rho}(\hat{x},{s}(\hat{x},\mu)+\tau\widetilde{\Delta s};\mu)\rangle\;d\tau$
$=-\int_{0}^{1}\int_{0}^{\tau}\langle\widetilde{\Delta s},D^{2}_{ss}\eta_{\rho}(\hat{x},{s}(\hat{x},\mu)+t\widetilde{\Delta s};\mu)\widetilde{\Delta s}\rangle\;dtd\tau$
${\leq\int_{0}^{1}\int_{0}^{\tau}\dfrac{\mu\widetilde{\delta}_{s,\rho}^{2}}{(1-\widetilde{\delta}_{s,\rho}+t\widetilde{\delta}_{s,\rho})^{2}}\;dtd\tau}$
$=\left(\frac{\tilde{\delta}_{s,\rho}}{1-\tilde{\delta}_{s,\rho}}+\ln(1-\tilde{\delta}_{s,\rho})\right){\mu}$
$\leq\frac{\kappa}{2}\mu,$
where the first inequality follows from Theorem 2.2 and Proposition 1. By the same argument, we conclude that $\theta_{\hat{x},\rho}(\mu)\leq\frac{\kappa}{2}\mu$. Therefore,
$\theta_{\rho}(\hat{x},s;\mu)=\theta_{s,\rho}(\mu)+\theta_{\hat{x},\rho}(\mu)\leq\kappa\mu.$ (121)
This completes the proof.
Remark 4
By (76) and Theorem 5.4, if $\xi_{\rho}(\hat{x},s;\mu)\leq\kappa$, then
$|\eta_{\rho}(\hat{x},s;\mu)-{\rm val}(\mathrm{P}_{\mu})|\leq\kappa\mu.$
Therefore, the PFSNM can be viewed as a relaxation of the IPM. The interpolation between the subproblem objectives of the two methods is controlled by the parameter $\mu$, which establishes a direct connection between the SNM and the IPM.
The following lemma relates the merit function $\xi_{\rho}(\hat{x},s;\mu)$ to the quantities $\nabla_{w}\eta_{\rho}(\hat{x},s;\mu)$ and $S_{\eta_{\rho}}(\hat{x},s;\mu)$, which is crucial for the subsequent complexity analysis.
Lemma 3
For any $\mu>0$, $\rho\geq 1$, and any point $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\xi_{\rho}(\hat{x},s;\mu)=\max\limits_{0\neq h\in\hat{\mathbb{E}}\times\mathbb{E}}\frac{\left|\langle\nabla_{w}\eta_{\rho}(\hat{x},s;\mu),h\rangle\right|}{\sqrt{\mu S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]}}.$ (122)
Proof
By Theorem 3.2, $\eta_{\rho}(\cdot,s;\mu)$ is nondegenerate $\mu$-self-concordant on $\hat{\mathbb{E}}$ for every $s\in\mathbb{E}$, and $\eta_{\rho}(\hat{x},\cdot;\mu)$ is nondegenerate $\mu$-self-concordant on $\mathbb{E}$ for every $\hat{x}\in\hat{\mathbb{E}}$. It follows from (Nesterov and Nemirovskii, 1994, Proposition 2.2.1) that for any nonzero direction $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\sqrt{\mu}\xi_{\hat{x},\rho}(\hat{x},s;\mu)\geq\frac{\left|\langle\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu),h_{\hat{x}}\rangle\right|}{\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)h_{\hat{x}}\rangle}}$, (123)
$\sqrt{\mu}\xi_{s,\rho}(\hat{x},s;\mu)\geq\frac{\left|\langle\nabla_{s}\eta_{\rho}(\hat{x},s;\mu),h_{s}\rangle\right|}{\sqrt{-\langle h_{s},D^{2}_{ss}\eta_{\rho}(\hat{x},s;\mu)h_{s}\rangle}}.$
Consequently, we have
$\sqrt{\mu}\xi_{\rho}\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}h_{\hat{x}}\rangle-\langle h_{s},D^{2}_{ss}\eta_{\rho}h_{s}\rangle}$ (124)
$=$ $\ \sqrt{\mu}\sqrt{\xi_{\hat{x},\rho}^{2}+\xi_{s,\rho}^{2}}\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}h_{\hat{x}}\rangle-\langle h_{s},D^{2}_{ss}\eta_{\rho}h_{s}\rangle}$
$\overset{(i)}{\geq}$ $\ \sqrt{\mu}\left(\xi_{\hat{x},\rho}\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}h_{\hat{x}}\rangle}+\xi_{s,\rho}\sqrt{-\langle h_{s},D^{2}_{ss}\eta_{\rho}h_{s}\rangle}\right)$
$\overset{(ii)}{\geq}$ $\left|\langle\nabla_{\hat{x}}\eta_{\rho},h_{\hat{x}}\rangle\right|+\left|\langle\nabla_{s}\eta_{\rho},h_{s}\rangle\right|$
$\geq$ $\left|\langle\nabla_{w}\eta_{\rho},h\rangle\right|,$
where both $(i)$ and $(ii)$ follow from the Cauchy–Schwarz inequality. Thus,
$\xi_{\rho}(\hat{x},s;\mu)\geq\frac{|\langle\nabla_{w}\eta_{\rho}(\hat{x},s;\mu),h\rangle|}{\sqrt{\mu S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]}},\quad\forall\,h\in\hat{\mathbb{E}}\times\mathbb{E},h\neq 0.$ (125)
Choosing
$h=\left((D_{\hat{x}\hat{x}}^{2}\eta_{\rho})^{-1}\nabla_{\hat{x}}\eta_{\rho},(D_{ss}^{2}\eta_{\rho})^{-1}\nabla_{s}\eta_{\rho}\right),$ (126)
inequality (125) holds with equality, which completes the proof.
Building on these estimates, we establish the main result of the paper: Algorithm 2 admits a polynomial-time complexity bound of order $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, matching the best‐known complexity of the classical short-step IPMs.
Theorem 5.5
Let $\rho\geq 1$ and choose
$1>\sigma\geq 1-\dfrac{\ln(\gamma)}{2\rho\sqrt{\nu}+\ln(\gamma)},\;where\;\gamma:=\dfrac{2\kappa+\sqrt{2}/\rho}{\kappa+\sqrt{2}/\rho}.$ (127)
Then Algorithm 2 requires at most $\mathcal{O}(\sqrt{\nu}\ln(\mu^{(0)}/\varepsilon))$ iterations to attain the desired accuracy $\varepsilon$.
Proof
First, we estimate the number of inner iterations, denoted by $N_{\text{in}}$. Fix an outer iteration index $k$ and suppose that $\xi_{\rho}(\hat{x}^{(k)},s^{(k)};\mu^{(k)})\leq\kappa$. For a fixed point $w^{(k)}=(\hat{x}^{(k)},s^{(k)})$ and $\rho\geq 1$, define the merit function $\psi:\mathbb{R}\times\hat{\mathbb{E}}\times\mathbb{E}\to\mathbb{R}$ as
$\psi(\mu,h):=\frac{\langle\nabla_{w}\eta_{\rho}(w^{(k)};\mu),h\rangle^{2}}{\mu S_{\eta_{\rho}}(w^{(k)};\mu)[h,h]}.$ (128)
A direct computation gives
$\psi^{\prime}(\mu,h)$ $=\frac{2\langle\nabla_{w}\eta_{\rho}(w^{(k)};\mu),h\rangle\cdot\langle\nabla_{w}\eta_{\rho}^{\prime}(w^{(k)};\mu),h\rangle}{\mu S_{\eta_{\rho}}(w^{(k)};\mu)[h,h]}-\frac{\langle\nabla_{w}\eta_{\rho}(w^{(k)};\mu),h\rangle^{2}}{\mu^{2}S_{\eta_{\rho}}(w^{(k)};\mu)[h,h]}$ (129)
$\qquad\qquad-\frac{\langle\nabla_{w}\eta_{\rho}(w^{(k)};\mu),h\rangle^{2}S^{\prime}_{\eta_{\rho}}(w^{(k)};\mu)[h,h]}{\mu(S_{\eta_{\rho}}(w^{(k)};\mu)[h,h])^{2}}.$
Let $\mu\in[\mu^{(k+1)},\mu^{(k)}]$. It follows from Theorems 5.25.3 that
$|\psi^{\prime}(\mu,h)|$ $\leq\sqrt{\frac{8\nu}{\mu^{2}}}\sqrt{\psi(\mu,h)}+\frac{1+\rho(1+2\sqrt{\nu})}{\mu}\psi(\mu,h)$ (130)
$\leq\sqrt{\frac{8\nu}{(\mu^{(k+1)})^{2}}}\sqrt{\psi(\mu,h)}+\frac{1+\rho(1+2\sqrt{\nu})}{\mu^{(k+1)}}\psi(\mu,h).$
Define $c_{1}:=\frac{1+\rho(1+2\sqrt{\nu})}{2\mu^{(k+1)}}$ and $\Psi(\mu,h):=e^{c_{1}\mu}\sqrt{\psi(\mu,h)}$. We have
$\Psi^{\prime}(\mu,h)\leq e^{c_{1}\mu}\frac{\sqrt{2\nu}}{\mu^{(k+1)}}.$ (131)
Integrating both sides of (131) over $\mu^{(k+1)},\mu^{(k)}]$ yields
$\Psi(\mu^{(k+1)},h)-\Psi(\mu^{(k)},h)\leq\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}\left(e^{c_{1}\mu^{(k)}}-e^{c_{1}\mu^{(k+1)}}\right).$ (132)
This implies that for any nonzero direction $h\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\sqrt{\psi(\mu^{(k+1)},h)}$ (133)
$\leq$ $\,e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}\sqrt{\psi(\mu^{(k)},h)}+\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}\left(e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}-1\right)$
$\leq$ $\,e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}\xi_{\rho}(\hat{x},s;\mu^{(k)})+\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}\left(e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}-1\right),$
where the last inequality follows from Lemma 3. Since (133) holds for every nonzero direction $h\in\hat{\mathbb{E}}\times\mathbb{E}$, we conclude by Lemma 3 again that
$\xi_{\rho}(\hat{x},s;\mu^{(k+1)})$ $\leq e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}\xi_{\rho}(\hat{x},s;\mu^{(k)})$ (134)
$\qquad\qquad\qquad+\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}\left(e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}-1\right).$
Recall $\mu^{(k+1)}=\sigma\mu^{(k)}$. Let
$\left\{\begin{array}[]{ll}&\alpha_{1}(\sigma,\rho):=e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}=e^{\frac{\rho(1+2\sqrt{\nu})+1}{2}(\frac{1}{\sigma}-1)},\\ &\alpha_{2}(\rho):=\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}=\frac{2\sqrt{2\nu}}{\rho(1+2\sqrt{\nu})+1}.\end{array}\right$. (135)
It can be verified that
$\alpha_{1}(\sigma,\rho)\leq e^{\frac{\rho(1+2\sqrt{\nu})+1}{2}\cdot\frac{\ln(\gamma)}{2\rho\sqrt{\nu}}}\leq\gamma\text{ \, and\, }\alpha_{2}(\rho)\leq\sqrt{2}/\rho,$ (136)
whenever $\rho\geq 1$ and $1>\sigma\geq 1-\dfrac{\ln(\gamma)}{2\rho\sqrt{\nu}+\ln(\gamma)}$. At the end of $k$-th iteration, one has $\xi_{\rho}(w^{(k)};\mu^{(k)})\leq\kappa$. It follows immediately that
$\xi_{\rho}(w^{(k)};\mu^{(k+1)})\leq\alpha_{1}(\sigma,\rho)\kappa+\alpha_{2}(\rho)(\alpha_{1}(\sigma,\rho)-1)\leq 2\kappa.$ (137)
By Theorem 2.3(iii), one Newton step suffices to obtain
$\xi_{\rho}(\hat{x}^{(k+1)},s^{(k+1)};\mu^{(k+1)})\leq\kappa.$
Hence, $N_{\rm in}=1$.
Next, we estimate the number of outer iterations, denoted by $N_{\text{out}}$. Since $\mu^{(k)}=\sigma^{k}\mu^{(0)}$, termination of Algorithm 2 at $k$-th iteration implies
$\mu^{(k)}\leq\varepsilon\quad\Longrightarrow\quad k\geq\frac{\ln({\mu^{(0)}}/{\varepsilon})}{\ln(1/\sigma)}.$ (138)
Consequently,
$N_{\text{out}}\leq-\ln\!\left(\frac{\mu^{(0)}}{\varepsilon}\right)/\ln(\sigma)+1=\mathcal{O}\!\left(\sqrt{\nu}\ln\!\left(\frac{\mu^{(0)}}{\varepsilon}\right)\right).$ (139)
The total number of iterations is given by
$N_{\text{in}}\times N_{\text{out}}=\mathcal{O}\left(\sqrt{\nu}\ln\left(\dfrac{\mu^{(0)}}{\varepsilon}\right)\right),$
which completes the proof.
Remark 5
Combining Theorems 5.1 and 5.5, we obtain that the first-phase of PFSNM requires at most $\mathcal{O}\left(\ln\left(1+\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}\right)\right)$ iterations, whereas the second-phase of PFSNM admits an iteration complexity of $\mathcal{O}\left(\sqrt{\nu}\ln\left(\mu^{(0)}/{\varepsilon}\right)\right)$. Therefore, the overall iteration complexity is
$\mathcal{O}\left(\ln\left(1+\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}\right)\right)+\mathcal{O}\left(\sqrt{\nu}\ln\left(\frac{\mu^{(0)}}{\varepsilon}\right)\right).$
Since $\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}$ is independent of $\varepsilon$, the above bound can be written more compactly as $\mathcal{O}\left(\sqrt{\nu}\ln\left(1/{\varepsilon}\right)\right)$.
6 Computational results
To validate the effectiveness of the proposed method (PFSNM), this section reports numerical results on three benchmarks and compares PFSNM with several widely used conic programming solvers, including SDPT3 Tütüncü et al. (2003), SeDuMi Sturm (1999), ECOS Domahidi et al. (2013), and Clarabel Goulart and Chen (2024). The test instances consist of linear programs from the NETLIB collection [^1] https://netlib.org/lp/data/, convex quadratic programs (QP) from the Maros–Mészáros collection [^2] https://www.doc.ic.ac.uk/~im/, and second-order cone programs arising from square-root Lasso formulations constructed using matrices from the SuiteSparse Matrix Collection [^3] https://sparse.tamu.edu/. All computational results are obtained on a Windows 10 personal computer equipped with an Intel i5-8300H processor (4 cores, 8 threads, 2.3 GHz) and 16 GB of RAM. The proposed method is implemented in C.
We evaluate solver performance employing performance profiles Dolan and Moré (2002) and the shifted geometric mean [^4] https://plato.asu.edu/ftp/shgeom.html (SGM). Let $\mathcal{P}$ denote the benchmark set and $\mathcal{S}$ the set of solvers. For each problem $p\in\mathcal{P}$ and solver $s\in\mathcal{S}$, let $t_{p,s}$ be the runtime, and define the performance ratio
$r_{p,s}=\frac{t_{p,s}}{\min_{s^{\prime}\in\mathcal{S}}t_{p,s^{\prime}}}\in[1,\infty],$ (140)
where $r_{p,s}=\infty$ if solver $s$ fails to solve problem $p$ within the time limit of 1000 seconds. The performance profile is given by
$\rho_{s}(\tau)=\frac{1}{|\mathcal{P}|}\Big|\big\{p\in\mathcal{P}:\ r_{p,s}\leq\tau\big\}\Big|,\qquad\tau\geq 1,$ (141)
which measures the fraction of instances for which $s$ is within a factor $\tau$ of the best solver. The value at $\tau=1$ reflects the frequency with which solver $s$ is the fastest, while the limiting value as $\tau$ grows measures its empirical success rate. In addition, we summarize the overall performance via the shifted geometric mean (with offset =1=1)
$\mathrm{SGM}_{s}=\exp\!\left(\sum_{p\in\mathcal{P}}\ln\left(\frac{\max\{1,t_{p,s}+\text{offset}\}}{|\mathcal{P}|}\right)\right)-\text{offset},$ (142)
so that smaller values indicate better aggregate efficiency while remaining insensitive to a small number of difficult instances.
6.1 Linear programs
We first test the solvers on linear programs from the widely used NETLIB collection. Each solver is run with the same time limit and accuracy requirements, and the results are summarized by Table 2 and Fig. 3.
Table 2 reports the solved-problem ratio together with the shifted geometric mean. The main observation is that PFSNM achieves the strongest overall combination of robustness and efficiency on this benchmark. In particular, it achieves the best aggregate runtime behavior according to the SGM, and it does so without compromising reliability, whereas competing solvers exhibit either a lower success rate or a noticeably larger SGM. This indicates that the advantage of PFSNM on NETLIB is consistent across instances, reflecting a favorable overall balance between convergence robustness and computational cost.
Table 2: SGMs of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the NETLIB collection.
Solver PFSNM SDPT3 SeDuMi ECOS Clarabel
Solved problems 100% 73.47% 93.88% 95.92% 97.96%
SGM 1.0000 25.0299 3.5985 2.7778 2.0520
Fig. 3 provides a more intuitive perspective through performance profiles. The curve of PFSNM stays above the competing solvers over essentially the entire range of $\tau$. This behavior indicates that PFSNM is frequently the fastest solver among the five solvers on a large portion of the benchmark set. Furthermore, $\rho_{s}(\tau)$ for PFSNM approaches 11 as $\tau$ increases, which means that it successfully solves all LP instances under the imposed limits. In contrast, the profiles of the other solvers level off below 11, and their slower rise for small $\tau$ indicates weaker competitiveness on instances with comparable runtimes.
Figure 3: Performance profiles of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the NETLIB collection.
6.2 Quadratic programs
We next consider convex quadratic programs from the Maros–Mészáros collection, a standard benchmark for QP problems. The results are summarized in Table 3 and Fig. 4.
Table 3 suggests that PFSNM is competitive in terms of aggregate efficiency, although it is not the top-performing solver overall. In particular, its shifted geometric mean is the second-best among the tested solvers (about $2.66$), whereas Clarabel attains the best value (normalized to $1.00$). At the same time, the solved-problem ratios indicate that Clarabel succeeds on a larger portion of the QP instances (about $91.3\%$), while PFSNM solves a smaller but still sizable fraction (about $82.6\%$). Overall, the table indicates that the main difference between PFSNM and the best solver on this benchmark lies in robustness on a subset of instances.
Table 3: SGMs of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the Maros–Mészáros collection.
Solver PFSNM SDPT3 SeDuMi ECOS Clarabel
Solved problems 82.61% 77.54% 83.33% 72.46% 91.30%
SGM 2.6598 8.1669 5.8094 7.7222 1.0000
Fig. 4 provides a consistent view. The performance profile of PFSNM rises quickly for small values of $\tau$, which indicates that when PFSNM succeeds, it often achieves runtimes close to the best solver on those instances. However, the limiting value of its profile remains below that of the most reliable solver, reflecting the gap in solved ratios reported in Table 3. Overall, the QP results show that PFSNM can be fast on the instances it solves, while improving robustness on the harder subset of Maros–Mészáros problems would further enhance its performance on this benchmark.
Figure 4: Performance profiles of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the Maros–Mészáros collection.
6.3 Second-order cone programs
Finally, we evaluate the solvers on a family of SOCP instances constructed from Lasso formulations. The data matrices are drawn from the SuiteSparse Matrix Collection, and each matrix is used to build a square-root Lasso problem Belloni et al. (2011); Liang et al. (2021) of the form
$\min\limits_{y\in\mathbb{R}^{n}}\left\{\|Dy-b\|_{2}+\varrho\|y\|_{1}\right\},$ (143)
where $D\in\mathbb{R}^{d\times n}$ is a matrix from the SuiteSparse Matrix Collection, $b\in\mathbb{R}^{d}$ is a given vector, and $\varrho$ is a penalty parameter. This problem is equivalent to the following SOCP instance:
$\min\left\{t+\varrho\sum\limits_{i=1}^{n}(y^{+}_{i}+y^{-}_{i})\,\Big|\,Dy^{+}-Dy^{-}-d=b,\,(t,d)\in\mathbb{Q}^{d+1},\,y^{+},y^{-}\in\mathbb{R}^{n}_{+}\right\}$ (144)
We follow the recent work Goulart and Chen (2024) and choose $\varrho=\|D^{\top}b\|_{\infty}$. The vector $b$ is set to be the all-ones vector. The results are reported in Table 4 and Fig. 5.
Table 4 shows that all solvers solve all the instances, so the comparison is driven by efficiency rather than robustness. Under this setting, PFSNM attains the smallest SGM (normalized to $1.0000$). The closest competitor is Clarabel, with a SGM only slightly larger (about $1.08$), whereas the remaining solvers have clearly larger SGM values. Hence, the table indicates that PFSNM provides the best aggregate runtime performance on these Lasso-type SOCPs, with a particularly tight competition against Clarabel.
Table 4: SGMs of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on SOCP problems constructed from SuiteSparse matrices.
Solver PFSNM SDPT3 SeDuMi ECOS Clarabel
Solved problems 100% 100% 100% 100% 100%
SGM 1.0000 3.5750 2.5645 2.3100 1.0818
Fig. 5 corroborates the table-based summary. Since all solvers succeed, the key difference lies in how quickly each profile rises near $\tau=1$. The PFSNM curve increases the fastest and stays close to the best observed curve over a wide range of $\tau$, meaning that it achieves the best or near-best runtime on a large fraction of instances. Combined with the SGM results, these experiments indicate that PFSNM offers strong and reliable performance on SOCP problems constructed from SuiteSparse matrices.
Figure 5: Performance profiles of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on SOCP problems constructed from SuiteSparse matrices.
7 Conclusion
The PFSNM has been proposed for SCP based on a reduced SBAL function. Its associated parameterized smooth system has been shown to be equivalent to the first-order optimality conditions of a structured minimax problem. This characterization makes it possible to analyze the method within a self-concordant convex-concave framework adapted to the reduced formulation. It has been proved that the reduced SBAL function is $\mu$-self-concordant convex-concave and that the resulting method attains a worst-case iteration complexity of $O(\sqrt{\nu}\ln(1/\varepsilon))$. This iteration complexity matches the best-known short-step bound for IPMs on symmetric cones. Moreover, the reduced formulation also yields Newton systems with an explicit Schur complement, which lowers the cost of system formation relative to existing smoothing Newton methods. Numerical results indicate that the method is competitive on standard conic benchmarks.
Appendix A Auxiliary proofs
A.1 Proof of Proposition 1
Proof
For any point $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and any direction $h_{\hat{x}}\in\hat{\mathbb{E}}$, define $h=(h_{\hat{x}},0)\in\mathbb{E}\times\mathbb{E}$ and let $\varrho(t)=D^{2}f(w+th)[h,h]=D^{2}_{\hat{x}\hat{x}}f(w+th)[h_{\hat{x}},h_{\hat{x}}$]. By the $\alpha$-self-concordant convex-concave property of $f$, we have
$\varrho^{\prime}(t)$ $=D^{3}_{\hat{x}\hat{x}\hat{x}}f(w+th)[h_{\hat{x}},h_{\hat{x}},h_{\hat{x}}$]
$=D^{3}f(w+th)[h,h,h]$
$\leq\frac{2}{\alpha^{1/2}}\left(S_{f}(w+th)[h,h]\right)^{3/2}$
$=\frac{2}{\alpha^{1/2}}\left(D^{2}_{\hat{x}\hat{x}}f(w+th)[h_{\hat{x}},h_{\hat{x}}]\right)^{3/2}.$
Taking $t=0$ yields
$D^{3}_{\hat{x}\hat{x}\hat{x}}f(\hat{x},s)[h_{\hat{x}},h_{\hat{x}},h_{\hat{x}}]\leq\frac{2}{\alpha^{1/2}}\left(D^{2}_{\hat{x}\hat{x}}f(\hat{x},s)[h_{\hat{x}},h_{\hat{x}}]\right)^{3/2}.$
[…]
The conclusion in $(ii)$ follows directly from (Nesterov and Nemirovskii, 1994, Proposition 9.1.1).
A.2 Proof of Theorem 2.3(v)
Proof
Define
$\xi_{\hat{x}}(w)=\Big(\tfrac{1}{\alpha}\big\langle\nabla_{\hat{x}}f(w),\big(D^{2}_{\hat{x}\hat{x}}f(w)\big)^{-1}\nabla_{\hat{x}}f(w)\big\rangle\Big)^{1/2},$ (145)
$\xi_{s}(w)=\Big(\tfrac{1}{\alpha}\big\langle\nabla_{s}f(w),\big(-D^{2}_{ss}f(w)\big)^{-1}\nabla_{s}f(w)\big\rangle\Big)^{1/2}.$
[…]
$\tilde{\delta}_{\hat{x}}(w)=\Big(\tfrac{1}{\alpha}\langle d,D^{2}_{\hat{x}\hat{x}}f(\hat{x},s)\,d\rangle\Big)^{1/2}.$ (146)
By Proposition 1, $\varrho_{s}(\cdot)$ is $\alpha$-self-concordant convex on $\hat{\mathbb{E}}$. Consequently, it follows from (Nesterov and Nemirovskii, 1994, Eq. (2.2.31)) that
$\tilde{\delta}_{\hat{x}}(w)\leq 1-(1-3\xi_{\hat{x}}(w))^{1/3}\leq 1-(1-3\xi(w))^{1/3}.$ (147)
Similarly, one has
$\tilde{\delta}_{s}(w)\leq 1-(1-3\xi_{s}(w))^{1/3}\leq 1-(1-3\xi(w))^{1/3}.$ (148)
Consequently
$\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}\leq 1-(1-3\xi(w))^{1/3}.$ (149)
Further, if $\xi(w)\leq 0.1$, then
$\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}\leq 1-0.7^{1/3}<0.2,$ (150)
which completes the proof.